### v1

# RAP-SAM [Uncaptioned image]: Towards Real-Time All-Purpose Segment Anything

Abstract
Advanced by transformer architecture, vision foundation models (VFMs) achieve remarkable progress in performance and generalization ability. Segment Anything Model (SAM) is one remarkable model that can achieve generalized segmentation. However, most VFMs cannot run in real-time, which makes it difficult to transfer them into several products. On the other hand, current real-time segmentation mainly has one purpose, such as semantic segmentation on the driving scene. We argue that diverse outputs are needed for real applications. Thus, this work explores a new real-time segmentation setting, named all-purpose segmentation in real-time, to transfer VFMs in real-time deployment. It contains three different tasks, including interactive segmentation, panoptic segmentation, and video segmentation. We aim to use one model to achieve the above tasks in real-time. We first benchmark several strong baselines. Then, we present Real-Time All Purpose SAM (RAP-SAM). It contains an efficient encoder and an efficient decoupled decoder to perform prompt-driven decoding. Moreover, we further explore different training strategies and tuning methods to boost co-training performance further. Our code and model are available at https://github.com/xushilin1/RAP-SAM/.
1 Introduction
Figure 1: We present real-time all-purpose segmentation to segment and recognize objects for image, video, and interactive inputs. In addition to benchmarking, we also propose a simple yet effective baseline, named RAP-SAM, which achieves the best accuracy and speed trade-off among three different tasks. Both real-time panoptic segmentation and video instance segmentation are shown at the bottom.
Recent advancements in computer vision, propelled by transformer architectures [13, 2, 75], focus on developing a singular, large-scale model versatile enough for various tasks, including detection and segmentation, with a general-purpose design. Specifically, several studies [74, 80] adopting identical architectures have demonstrated superior performance compared to models designed exclusively for specific tasks. Furthermore, numerous studies [34, 30] investigate the foundation model design using extensive data. A notable example, the Segment Anything Model (SAM) [34], introduces an interactive model driven by visual prompts for versatile segmentation. Concurrently, for semantic-level segmentation, numerous studies [10, 93, 47, 46, 29] employing a universal design have surpassed previous models tailored for specific image and video segmentation tasks.
However, the majority of these studies face challenges in achieving real-time performance or compatibility with mobile devices, primarily due to their heavy encoders and cascaded decoders. This constraint significantly limits the applicability of these advanced architectures in practical vision applications, although they can achieve high performance. On the contrary, various studies explore real-time segmentation using distinctive designs. Yet, these efforts predominantly concentrate on singular application purposes, including autonomous driving scenes, the COCO dataset for panoptic segmentation (PS) [50], and real-time video instance/object segmentation (VIS) [81]. Currently, no studies investigate real-time, all-purpose segmentation, wherein a model is capable of performing universal segmentation tasks, encompassing image segmentation, video instance/video segmentation, and SAM-like interactive segmentation.
Inspired by the aforementioned challenges, this study investigates the novel issue of real-time, all-purpose segmentation. We pose a critical inquiry: given the limited computational resources and model capacity, how can we develop an efficient, all-purpose segmentation model? This entails creating a single model capable of segmenting, tracking, and classifying each pixel in real-time, akin to performing interactive segmentation similar to SAM.
Initially, we define this research problem by integrating existing benchmarks. Specifically, we use the COCO [50] and YouTube-VIS 2019 [81] datasets for joint co-training, employing identical hyper-parameters. Beyond the semantic queries used for semantic-level segmentation in prior research, our approach additionally incorporates visual prompt queries in SAM for interactive segmentation. However, there are still several challenges: 1), No previous works explore joint co-training for three different purposes. Thus, extensive experiments are needed to search for a better meta-architecture. 2), Interactive segmentation and semantic-level segmentation need better strategies to balance each other. 3), Under the real-time setting, several previous architectures [80, 42] for multi-task segmentation may not be practical.
To address these challenges, we benchmark several recent works [85, 10, 25] in a real-time setting. Then, we initially investigate a range of meta-architectures suitable for all-purpose segmentation. Unlike previous works, we avoid cascaded architecture and only use the pyramid feature once when performing dynamic convolution. We explore four different architectures in the case of various decoder designs, where we adopt a single decoder with multiple queries. Via extensive experiments, we adopt pooling-based dynamic convolution to replace per-pixel cross-attention to achieve better accuracy and speed trade-off. To balance interactive and semantic-level segmentation, we present an effective dual adapter design that better adapts shared knowledge of the same decoder. Finally, we term our best model as Real-time All-Purpose SAM (RAP-SAM). To the best of our knowledge, it is the first real-time all-purpose segmentation model. Moreover, due to the semantic level and interactive level segmentation ability, our model can also create more new applications, such as interactive video segmentation. Extensive experiments show the effectiveness of our design. The main contributions of this work are:
  • We introduce all-purpose segmentation, a multi-task segmentation that aims to segment objects for image, video, and interactive inputs in real-time.
  • We benchmark several real-time transformer-based segmentation approaches for the new settings.
  • We present a simple yet fast baseline named RAP-SAM. It contains a lightweight feature extractor, a unified decoder, and two asymmetric adapters.
  • Extensive experiments demonstrate that RAP-SAM achieves the best speed and accuracy trade-off in the proposed benchmark and regular real-time semantic and panoptic segmentation benchmarks. We also show scalability across datasets and application demos.
2 Related Work
Universal Segmentation. Prior research has focused on designing segmentation models tailored for specific tasks. Recent developments in this field [13, 52, 2, 93, 10, 85, 49, 39, 17, 86, 42, 43, 45, 87] have shifted towards employing mask classification architectures, coupled with an end-to-end set prediction objective, to facilitate universal segmentation. This approach has outperformed specialized models [4, 33, 20, 48, 27, 88, 96, 7, 40, 41, 6] across various segmentation tasks, including those involving images, videos, and point clouds [32, 47, 46, 9, 67, 78, 79]. Specifically, Mask2Former [10] uses a masked-attention mechanism and surpasses the performance of prior specific image segmentation models. Similarly, Tube-Link [46] outperforms earlier models [47, 40, 51] tailored for specific video segmentation tasks. Semantic-SAM [35], built on Mask-DINO [36], employs a hybrid query design to facilitate both panoptic and interactive segmentation within a single framework. However, despite the strong performance of these works, their complexity, stemming from the heavier encoder and the cascaded decoder designs, poses significant challenges for integration into actual products. These are primarily hindered by slow processing speeds and substantial parameter overhead.
Vision Foundation Model. Recent advancements in the vision research community have been propelled by the development of Vision Foundation Models (VFMs), including pure-visual pre-training [19], vision-language pre-training [15, 37], and multi-modal models employing visual prompting [75, 1]. SAM [34] stands as a seminal work targeting general interactive segmentation. Subsequent studies [3, 12, 77] have adapted SAM for diverse applications, predominantly employing it as a segmentation tool. However, these Vision Foundation Models typically employ substantial visual encoders, like ViT-base [13], which poses challenges for deployment on real-time or mobile devices due to their size.
Efficient Model Design. This research direction primarily concentrates on the development of efficient CNNs [24, 66, 23, 28, 94, 56, 18], transformers [57, 64, 58], and hybrid architectures [90, 55, 8, 38, 100], aimed at advancing visual representation learning. Most research investigates backbone design within the constraints of real-time processing and parameter efficiency. These studies are orthogonal to our own, which primarily explores the impact of data, task, and decoder design. Additionally, our work employs efficient models as encoders and presents detailed benchmarks and analyses of the encoder’s effects.
Efficient Segmentation. Previous studies [95, 44, 84, 25, 22, 71, 83, 59] on efficient segmentation have predominantly concentrated on closed-set and specific domains. In particular, a significant portion of this research [44, 25, 60] is dedicated to driving scenarios. Recently, various studies [92, 71, 99] have developed efficient segmentation techniques that facilitate model execution on mobile devices. Mobile SAM [89] introduces a streamlined encoder distillation method. Fast SAM [97] employs a single-stage instance segmentation framework that directly decodes class-agnostic masks. Recently, multiple studies have been conducted on efficient panoptic segmentation [69, 25, 63] and rapid video instance segmentation [82, 91]. However, these real-time segmentation methods are limited to specific tasks. We contend that an all-purpose model capable of real-time performance would have wide-ranging applications in editing, tracking, and segmentation functionalities within various products.
3 Problem Setup
Settings. Our proposed real-time all-purpose segmentation contains three visual inputs, including image, video, and visual prompts (boxes and points). It outputs corresponding masks, labels, and instance ID. We adopt panoptic segmentation for image segmentation, which unifies semantic and instance segmentation. For video segmentation, we adopt video instance segmentation, where we aim to segment and track each foreground instance. For interactive segmentation, we follow the SAM-like setting, where we perform class-agnostic segmentation according to the user’s inputs. Compared with previous settings, as shown in Tab. 1, we can achieve more flexible input and output, which is more challenging. Combined with interactive segmentation and video segmentation, we can also achieve interactive video segmentation as video object segmentation.
Datasets. For image segmentation and interactive segmentation, we use the well-known COCO dataset [50] for benchmarking. We adopt partially SAM data [34] for testing. Due to the limited resources and limited model capacity in real-time settings, we do not use SAM data for training. We take COCO data as SAM data for training and testing, where the boxes and random points sampled from masks are used for visual prompting. For video segmentation, we adopt the widely used YouTube-VIS 2019 dataset [81] for training. We do not use the SAM data for our benchmarking. There are several reasons: (1), SAM data is not well annotated as the COCO dataset. (2), We only focus on object-level and scene-level segmentation, while SAM data contain more granularity. (3), SAM data has no semantics, while our goal is to predict mask labels for real applications.
Figure 2: RAP SAM overview. Our method contains three visual inputs: image, video, and visual prompts. Utilizing positional encoding, we generate prompt queries from these visual prompts. The learnable object queries, alongside the prompt queries and the feature map $F$, are directed to the multi-stage decoder. This process generates multi-stage predictions and refined queries. These refined queries engage in cross-attention with $F$, resulting in the final prediction.
4 Proposed Method
Motivation. Employing the same formulation, we first revisit three tasks: panoptic segmentation, video instance segmentation, and interactive segmentation. When combined with a query-based mask transformer, object queries can represent all entities within the scene, encompassing image-level objects, video-level tracked objects, and prompt-based specific objects. Each task’s formulation is as follows:
For panoptic segmentation [33], given an input image $I\in\mathbb{R}^{H\times{W}\times 3}$, the objective is to generate a group of masks $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i})\}_{i=1}^{N}\,$ where $c_{i}$ denotes the ground truth class label of the binary mask $m_{i}$ and $N$ is the number of masks, ${H}\times{W}$ indicates the spatial dimensions.
For video instance segmentation [81], given a video clip input $V\in\mathbb{R}^{T\times H\times{W}\times 3}$, where $T$ denotes the number of frames, the process aims to generate an instance mask tube $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i},d_{i})\}_{i=1}^{N}\,$, where $N$ represents the number of the tube masks $m_{i}\in{\{0,1\}}^{{T}\times{H}\times{W}}$, $c_{i}$ indicates the class label of the tube $m_{i}$, and $d_{i}$ identifies the instance ID for each tube.
Table 1: Real-Time Segmentation Settings. Our all-purpose segmentation supports more tasks in one framework.
Property Image Seg Video Seg SAM-Like Ours
Image Masks
Video Masks
Interactive
Semantic Labels
Multi tasks
Table 2: Comparison of Segmentation Methods. Our proposed RAP-SAM supports various segmentation tasks, and it can run in real time.
Methods SS PS VIS Interactive Multi-Task in One Model Real Time
ICNet [95]
Bi-Seg [84]
YOSO [25]
Mobilie-VIS [91]
SAM [34]
Mask2Former [10]
Video K-Net [47]
OneFormer [29]
RAP-SAM (Ours)
For SAM-like interactive segmentation [34], both the image $I$ and visual prompts $P\in\mathbb{R}^{K\times 4}$ (box prompts are used here) are taken as inputs. These prompts, which include points and boxes, lead to the outputs of corresponding binary image masks $\{y_{i}\}_{i=1}^{K}=\{m_{i}\in\{0,1\}^{H\times W}\}_{i=1}^{K}$, where $K$ denotes the number of visual prompts. Each visual prompt is encoded into one object query, which naturally can be the input of the decoder, as implemented in  [34, 10].
Single Object Query For All Segmentation. As mentioned above, by combining all three settings into one, we can represent all the output segmentation entities using the same query-based mask classification framework. In particular, one object query corresponds to one mask $m_{i}$, one label $c_{i}$, and one ID $d_{i}$. We aim to design a simple, efficient, and parameter-sharing framework to support these tasks.
Compared Baselines. In Sec. 5.1, we benchmark several baselines for our all-purpose segmentation. We use the prompt encoder (proposed in SAM [34]) to encode the visual prompt and jointly learn with object query for both PS and VIS on a shared decoder. Although we adopt the lightweight design to speed up, their performance is still limited or not balanced, as shown in Tab 3.
4.1 RAP-SAM Architecture
Overall Architecture. Our RAP-SAM is a simple encoder and decoder architecture. As shown in Fig. 2, it contains a backbone, a lightweight neck, and a shared multitask decoder. Visual prompts $P$ are also the input of the decoder. Following SAM, we also adopt the prompt encoder to encode visual prompts into a query. We adopt the same decoder for both visual prompts and initial object queries to share more computation and parameters. However, the goals of both are different. The former pays more attention to the local details, while the latter also considers the scene and temporal features. To better balance the results for interactive segmentation and image/video segmentation, we design a prompt adapter and an object adapter in the end of the decoder.
Lite Feature Extractor. Due to computation cost limitation, we avoid the design of previous methods, such as a large backbone and heavier transformer encoder. We explore lightweight backbones, including ResNet18 [21], STDC-v1 [14], and SeaFormer [71]. We adopt feature pyramid networks with deformable convolutions to fuse multi-scale features and obtain more aligned feature representation, motivated by previous works [44] on real-time semantic segmentation. For video input, we extract the spatial-temporal features. For simplicity, we use $F_{img}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times d}$ for image inputs and $F_{vid}\in\mathbb{R}^{T\times\frac{H}{4}\times\frac{W}{4}\times d}$ for video inputs.
Unified Dynamic Convolution Decoder. As in previous works, the goal of the decoder is to refine the object query. However, many of these approaches are impractical for real-time settings due to their reliance on heavily cascaded layers and pixel-wise cross-attention mechanisms. In contrast, our method employs a pooling-based dynamic convolution framework [93, 47] to enhance the decoder’s efficiency.
Given object query $Q_{i}\in\mathbb{R}^{N\times d}$ and prompt query $P_{i}\in\mathbb{R}^{K\times d}$, the initial masks $M_{i}$ are obtained by dot product with $F_{img}$ or $F_{vid}$. $i$ denotes the index of the decoder stage. According to the inputs, the masks can be image masks for panoptic segmentation $M_{i}^{pan}\in\mathbb{R}^{N\times\frac{H}{4}\times\frac{W}{4}}$ and interactive segmentation $M_{i}^{iter}\in\mathbb{R}^{K\times\frac{H}{4}\times\frac{W}{4}}$ or tube masks $M_{i}^{tube}\in\mathbb{R}^{N\times T\times\frac{H}{4}\times\frac{W}{4}}$. Then, we can obtain the corresponding query features $X_{i}^{pan}\in\mathbb{R}^{N\times d}$, $X_{i}^{iter}\in\mathbb{R}^{K\times d}$, $X_{i}^{tube}\in\mathbb{R}^{N\times d}$ via mask pooling:
$X_{i}^{pan}=\sum_{u}^{W}\sum_{v}^{H}M_{i-1}^{pan}(u,v)\cdot F_{img}(u,v).$ (1)
For video tasks, we group the spatial-temporal features as follows:
$X_{i}^{tube}=\sum_{t}^{T}\sum_{u}^{W}\sum_{v}^{H}M_{i-1}^{tube}(t,u,v)\cdot F_{vid}(t,u,v),$ (2)
where $u$, $v$ are the indices of spatial location, $t$ is index of input frame. Since the dynamic convolution performs similarly for each task, we use panoptic segmentation $X_{i}^{pan}$ as an illustration. For interactive segmentation, we adopt the same equation (Equ. 1). In our implementation, the learned query is shared for both $Q_{i}^{pan}$ and $Q_{i}^{tube}$.
In particular, we adopt the design [93, 68, 42] to refine input queries $Q_{i}^{pan}$ with features $X_{i}^{pan}$ which are grouped from their masks,
$\hat{Q}^{pan}_{i}=\mathrm{DynamicConv}(X_{i}^{pan},Q^{pan}_{i-1}),$ (3)
where the dynamic convolution uses the query features $X^{pan}_{i}$ to generate parameters to weight input queries $Q_{i-1}^{pan}$. Specifically, $\mathrm{DynamicConv}$ uses input query features $X^{pan}_{i}$ to generate gating parameters via a multilayer perceptron (MLP) and multiply back to the original query input $Q^{pan}_{i-1}$. We adopt the same design [93, 68] by learning gating functions to update the refined queries. The $\mathrm{DynamicConv}$ operation is based on:
$\hat{Q}^{pan}_{i}=\mathrm{Gate}_{x}(X^{pan}_{i})X^{pan}_{i}+\mathrm{Gate}_{q}(X^{pan}_{i})Q^{pan}_{i-1},$ (4)
where $\mathrm{Gate}$ is implemented with a fully connected (FC) layer followed by Layer Norm (LN) and a sigmoid layer. We adopt two different gate functions, including $\mathrm{Gate}_{x}$ and $\mathrm{Gate}_{q}$. They are implemented by MLP. The former is to weigh the query features, while the latter is to weigh corresponding queries. Next, we adopt one self-attention layer with feed-forward layers [70, 72] to learn the correspondence among each query and update them accordingly. This operation leads to the full correlation among queries, shown as follows:
$Q_{i}=\mathrm{FFN}(\mathrm{MHSA}(\hat{Q}_{i})+\hat{Q}_{i}),$ (5)
where $\mathrm{MHSA}$ means Multi Head Self Attention, $\mathrm{FFN}$ is the Feed Forward Network commonly used in the current version of Transformers [2, 13]. Finally, the refined masks are obtained via dot product between the refined queries $Q^{pan}_{i}$, and the input features $F_{p}$, $F_{s}$. For mask classification, we adopt several feed-forward layers on $Q^{pan}_{i}$ and directly output the class scores. Then, we perform the inner product between the learned queries $Q^{pan}_{i}$, $Q^{iter}_{i}$, and their features $F_{vid}$ and $F_{img}$. For the video segmentation task, the only difference is that we predict the tube mask rather than the image mask. The entire process (Equ. 1, Equ. 2, Equ. 3 and Equ. 5) is repeated three times in total.
Moreover, we also explore various decoupled designs, as shown in Fig. 3. In particular, as previous works suggest, we also explore decoupled decoders and their corresponding adapters. Via our detailed experiments, we do not find extra gains in the cases of decoupled decoders that introduce more computation and parameter costs. Please refer to Sec. 5.2 for the detailed decoder design.
Table 3: Real-Time All-Purpose Segmentation benchmark. Our proposed RAP-SAM achieves the best accuracy and speed trade-off on three tasks.
Method Backbone COCO-Panoptic COCO-SAM YouTube-VIS 2019 Flops Parameters FPS
PQ SQ $\mathrm{PQ_{th}}$ $\mathrm{PQ_{st}}$ mIoU mAP
Mask2Former [10] R18 35.6 77.5 39.0 30.3 54.7 38.6 89.8G 18.6M 31.2
Mask2Former [10] R50 42.9 79.8 47.6 35.6 58.0 42.1 153G 45.2M 26.6
MaskFormer [11] R18 31.0 76.0 33.2 27.7 55.6 34.1 79.7G 18.5M 38.0
MaskFormer [11] R50 37.4 77.3 41.2 31.8 58.8 40.0 143.0G 45.2M 34.3
MaskFormer [11] TopFormer [92] 31.6 75.9 33.5 28.6 56.1 38.2 60.0G 12.8M 29.9
kMaX-DeepLab [85] R18 27.8 71.0 30.3 24.1 16.9 23.1 87.1G 18.7M 15.0
kMaX-DeepLab [85] R50 36.9 75.7 42.7 28.4 20.1 26.4 280.0G 51.4M 14.9
YOSO [25] R18 31.6 76.8 36.4 24.4 45.7 31.4 57.3G 18.7M 41.0
YOSO [25] R50 37.3 76.9 42.7 29.2 49.2 40.7 119.0G 45.1M 36.3
YOSO [25] TopFormer [92] 31.0 75.9 33.0 26.9 45.1 27.9 36.8G 12.9M 31.1
RAP-SAM R18 39.9 78.6 43.3 34.8 52.7 38.7 60.5G 22.8M 40.3
RAP-SAM R50 46.9 80.8 51.6 39.8 57.9 46.2 123.0G 47.2M 35.1
RAP-SAM TopFormer [92] 34.6 77.0 37.8 29.8 53.3 41.7 40.1G 15.0M 30.7
Light-Weight Decoupled Adapter. After the shared decoder, we also add two lightweight adapters, $A_{obj}$ and $A_{prompt}$ to adapt the shared decoder’s knowledge better. In particular, we use the asymmetric design. $A_{obj}$ uses the same dynamic convolution design to refine the object query further, while $A_{prompt}$ uses pixel-wise cross-attention design. Our key insights are: (1) For panoptic segmentation and video segmentation, the interaction of the scene feature is necessary, but not for interactive segmentation. (2) despite the decoder being shared, the details are still missing for interactive segmentation due to the pooling effect in Equ. 1, where we even find inferior results after adopting a pooling-based adapter. More detailed adapter design can be found in Sec. 5.2.
4.2 Training and Inference
Joint Image and Video Segmentation Co-training. Our goal is to train all segmentation tasks only once jointly. All training targets are one entity label and mask for all three different cases. The entity can be thing, stuff, class-agnostic masks, and their corresponding labels. Note that the instance masks with the same ID $d$ form the tube masks. During training, we apply Hungarian matching between the predicted and ground-truth entity masks to assign object queries to video/image entities and then supervise their predicted masks and classification. The classifier is replaced by CLIP text embedding to avoid cross-dataset taxonomy conflicts.
Following previous works [10, 46], the final loss function is given as $L=\lambda_{cls}L_{m\_cls}+\lambda_{ce}L_{m\_ce}+\lambda_{dice}L_{m\_dice}$. Here, $L_{m\_cls}$ is the Cross-Entropy (CE) loss for mask classification, and $L_{m\_ce}$ and $L_{m\_dice}$ are mask Cross-Entropy (CE) loss and Dice loss [62, 73] for segmentation, respectively. By default, we set $\lambda_{cls}=2,\lambda_{ce}=5,\lambda_{dice}=5$.
Inference. we follow the same inference procedure of Mask2Former [10] for image segmentation. We merge the things and stuff for PS according to the sorted scores. The scores are generated by CLIP text embedding. For video instance segmentation, we use query matching rather than introducing extra tracking components to generate instance ID, following previous work [47, 26, 46]. We adopt a near-online method for inference. For interactive segmentation tasks, we follow the origin SAM [34], by providing box and point prompts and obtaining the binary masks. All the parameters are shared across three different tasks.
Extension and Application. Since our model can perform various segmentation tasks, RAP-SAM can be extended to more datasets and tasks, including semantic segmentation on ADE20k [98] or video panoptic segmentation on VIPSeg [61]. Moreover, our method can support prompt-driven video object segmentation by combining interactive segmentation with a video instance segmentation framework. We put the details of the extension and application for RAP-SAM in the supplementary.
5 Experiments
Datasets. Our benchmark uses COCO [50] and YouTube-VIS 2019 [81] datasets for benchmarking. COCO-SAM is constructed by using ground truth boxes and masks as visual prompt inputs, following the SAM-1B dataset [34]. In addition, to verify the effectiveness and generality of RAP-SAM, we also use other datasets, including ADE-20K [98] and VIP-Seg dataset [61] in Sec. 5.1.
Evaluation Metrics and Devices. For image segmentation, we adopt panoptic quality (PQ), which can be further decomposed into the segmentation quality (SQ) and the recognition quality (RQ). For the VIS task, mAP is adopted. We further report the semantic segmentation results (mIoU) and video panoptic segmentation results (STQ [76] and VPQ [31]). We adopt one A100 GPU for speed testing.
Re-implemented Baselines. Since no previous works explore joint co-training for image, video, and interactive segmentation in one framework. We re-implement several representative baselines using the same codebase, including non-real-time models (Mask2Former [10], kMaX-DeepLab [85]) for reference. In particular, we benchmark over ten different methods, including different backbones and decoders.
Implementation Details. We implement our models and all other baselines in PyTorch [65] with the MMDetection toolbox [5]. We use the distributed training framework with 16 A100 GPUs. Each mini-batch has two images per GPU. In particular, we adopt pseudo video training on COCO by moving image masks with random directions. All the models are trained with 24 epochs. For data augmentation, we adopt large-scale jitter as previous works [10] to build strong baselines. For all models, we adopt training steps and optimizers. Refer to the supplementary for more details.
5.1 Main Results
Our Benchmark Results. We list our benchmark results in Tab. 3. We benchmark recent state-of-the-art methods using the same training and test settings. From the table, our proposed RAP-SAM achieves the best speed and accuracy trade-off on all three different visual segmentation tasks, under the various backbones. Mask2Former [10] achieves similar or partially stronger results than our method. However, their speed is still limited, which is 7-8 FPS slower than our methods. YOSO [25] runs as fast as our method. However, the performance gap is still significant. Thus, our proposed RAP-SAM is a simple yet effective baseline for real-time and all-purpose settings.
Promptable Results Compared with SAM. In Tab. 4, we verify our trained model as a promptable segmenter, like SAM. We compare our method with SAM-Huge, which is much larger. Despite there being a gap, our method can run fast and with much less GFlops.
Comparison with Specific Design Models on VIP-Seg. We also verify the effectiveness of RAP-SAM on a more challenging video segmentation task, video panoptic segmentation on the VIP-Seg dataset. Compared with recent works [46, 61], RAP-SAM also archives the best speed and accuracy trade-off.
Table 4: Comparison with SAM [34] as a promptable segmenter on various detectors and instance segmenters.
Method Detectors mAP AP50 AP75 Parameters Flops
SAM-Huge Mask-RCNN (R50) 35.6 54.9 38.4 641M 3000G
Ours Mask-RCNN (R50) 26.7 43.0 27.9 47M 188G
SAM-Huge Mask-RCNN (X-101-64x4d) 39.0 60.5 42.1 641M 3000G
Ours Mask-RCNN (X-101-64x4d) 28.5 46.4 29.9 47M 188G
Figure 3: Meta-architecture exploration. (a), Simple shared decoder design. (b), Decoupled decoder design with two heads. (c), Shared decoder with decoupled adapter. (d), Decoupled decoder with the decoupled adapters. Best viewed in color and zoom in.
Table 5: Comparison with real-time video segmentation method on VIP-Seg validation set.
Method backbone $\rm VPQ^{1}$ $\rm VPQ^{2}$ $\rm VPQ^{4}$ $\rm VPQ^{6}$ VPQ STQ FPS
Clip-PanoFCN [61] R50 24.3 23.5 22.4 21.6 22.9 31.5 8
Video K-Net [47] R50 29.5 26.5 24.5 23.7 26.1 33.1 10
Tube-Link [46] STDCv1 32.1 31.3 30.1 29.1 30.6 32.0 14
Tube-Link [46] STDCv2 33.2 31.8 30.6 29.6 31.4 32.8 12
Ours R18 34.2 32.1 30.8 30.2 32.5 33.7 30
Table 6: Comparison with real-time panoptic segmentation method on ADE-20k validation set.
Method Backbone PQ $\mathrm{PQ^{th}}$ $\mathrm{PQ^{st}}$ Parameters Flops
PanSegFormer [49] R50 36.4 35.3 38.6 51M 214G
MaskFormer [11] R50 34.7 32.2 39.7 45.2M 181G
Mask2Former [10] R50 39.7 39.0 40.9 45.2M 226G
YOSO [25] R50 38.0 37.3 39.4 45.1M 176G
RAP-SAM R50 38.3 36.6 39.7 47.2M 179G
Comparison with Specific Design Models on ADE-20k. We further transfer RAP-SAM on ADE-20k datasets. Since our method is real-time model, we only list several representative works for reference. Again, compared with recent work [25], our method still achieves stronger results.
5.2 Ablation Study and Visual Analysis
Ablation on Shared Decoder Design. In Tab. 7, we find using simple pooling-based dynamic convolution performs well under real-time settings, where we adopt R18 as the backbone and keep the adapter unchanged. In particular, there is no difference in panoptic segmentation results and a slight gap for interactive segmentation. However, per-pixel cross-attention introduces extra computation costs.
Ablation on Meta-Architecture Design. In Tab. 8, we further explore the meta-architecture design, as shown in Fig. 3. We use R50 as the backbone and keep the adapter unchanged. From the table, we find using a shared decoder architecture (in Tab. 7) achieves the best parameter and performance trade-off. These findings may conflict with recent multi-task works [42] since the decoupled decoders may have more capacity. We argue that since our feature extractors are weaker than those works and the representation power is far from the heavier backbone such as Swin [52] or ViT-large [13], adding more capacity in the decoder may not boost the final performance.
Table 7: Ablation on shared decoder design.
Setting COCO-Panoptic COCO-SAM FPS
Per-Pixel Cross-Attention [10] 45.0 55.3 28.0
Pooling + Dynamic Convolution [68] 43.8 55.7 36.2
Pooling + Dynamic Convolution with Gate [93] 44.6 56.7 34.5
Table 8: Ablation on meta-architecture.
Setting COCO-Panoptic COCO-SAM Parameters
Fig. 3 (a) 43.8 54.2 46.3M
Fig. 3 (b) 44.0 55.2 53.6M
Fig. 3 (c) 44.6 56.7 47.3M
Fig. 3 (d) 45.2 56.2 54.6M
Table 9: Ablation on adapter design. CA: cross-attention. DC: dynamic convolution.
$\mathrm{A_{obj}}$ $\mathrm{A_{prompt}}$ COCO-Panoptic COCO-SAM
- - 44.2 53.2
CA CA 42.6 54.3
DC DC 44.7 52.1
DC CA 44.6 56.7
CA DC 44.5 56.2
Table 10: Ablation on joint co-training. (a), COCO-Panoptic. (b), YT-VIS 2019. (c), COCO-SAM.
Setting COCO-Panoptic YT-VIS 2019 COCO-SAM
a 36.6 - -
b - 21.5 -
a + b 36.2 36.0 50.7
a + b + c 35.7 35.3 50.9
Figure 4: The visualization results of our model on the YouTube-VIS 2019 dataset and the COCO dataset. The first three rows visualize five frames of inputs. The same instances are shown with the same color. The fourth row shows the interactive segmentation results with a single-point prompt (shown in green color). The last row shows the panoptic segmentation results.
Ablation on Adapter Design. Then, we explore the adapter design in Tab. 9. In particular, we use a pre-trained model without adapters for initialization to see the effect of different adapters. In particular, we find using asymmetric adapters works well for balanced results for object queries and prompt queries, since the goals of the two queries are different. The former needs the spatio-temporal and scene-level context, while the latter only cares about the region context with the input position as the guidance.
Effectiveness of Joint Co-training. We further explore the effect of each dataset in Tab. 10. We find joint co-training with image and video data leads to better performance for video instance segmentation. Adding COCO-SAM training leads to a few performance drops. This is because the learning targets are different for object queries and prompt queries.
Visualization. In Fig. 4, we give the three different visualization results on the YouTube-VIS 2019 dataset and the COCO dataset. We generate these visualization results using only a single model thanks to the joint co-training strategy. The first three rows in Fig. 4 show the video instance segmentation. The fourth row shows the interactive segmentation results with a single-point prompt (shown in green color), and the last row shows the panoptic results.
6 Conclusion
In this work, we explore a challenging real-time setting, all-purpose segmentation, to transfer more segmentation tasks into one model. To solve this problem, we introduce RAP-SAM, a real-time segmentation model, that can segment and track objects in image, video, and interactive modes. We design a simple shared decoder with two key designs: shared dynamic convolution and asymmetric adapters, which lead to the best trade-off between speed and accuracy on three segmentation tasks. We hope our proposed benchmark inspires future research on this new setting.
Overview. In this supplementary, we first present more detailed results to support both the benchmark and our proposed RAP-SAM in Sec. A. Then, in Sec. B, we present a more detailed visual comparison using our model. Finally, in Sec. C, we discuss several failure cases and potential future work.
Appendix A More Experiment Results
Benchmark More Methods. In Tab. 11, we benchmark more models and backbones [93, 57, 58] as the supplementary to main paper. Again, for various backbones and different heads, our model can achieve better or comparable results across three different tasks, while running in real-time. In particular, compared with the dynamic kernel-based method, K-Net, our approach achieves better performance on the R18 and R50 backbones and comparable results on the lightweight backbones.
Scale Up Joint Co-training. In the bottom of Tab. 11, we further scale up our model on a large backbone, ConvNeXt large [53]. From the table, we find there are significant gains in all settings while still running with 8 FPS. This means our proposed all-purposed segmentation still has room to balance the accuracy and inference speed.
Table 11: Benchmark More Models on Real-Time All-Purpose Segmentation. The GFlops are obtained with $1333\times 800$ inputs.
Method Backbone COCO-Panoptic COCO-SAM YouTube-VIS 2019 Flops Parameters FPS
PQ SQ $\mathrm{PQ_{th}}$ $\mathrm{PQ_{st}}$ mIoU mAP
YOSO [25] EdgeNexT [57] 31.8 74.5 35.6 26.1 47.4 35.4 40G 15.3M 28.6
K-Net [93] R18 [21] 33.1 75.7 36.8 27.4 53.6 34.8 124G 25.6M 30.4
K-Net [93] R50 [21] 39.2 78.2 43.9 31.6 56.3 41.7 171G 38.5M 20.9
K-Net [93] EdgeNexT [57] 38.0 76.7 42.5 31.4 55.7 44.0 108G 19.5M 30.5
Mask2Former [10] EdgeNexT [57] 39.8 78.8 43.6 34.0 54.2 - 73G 12.0M 25.6
RAP-SAM R18 39.9 78.6 43.3 34.8 52.7 38.7 60.5G 22.8M 40.3
RAP-SAM R50 46.9 80.8 51.6 39.8 57.9 46.2 123.0G 47.2M 35.1
RAP-SAM EdgeNexT [57] 38.0 77.7 42.1 31.9 54.8 46.3 44G 14.3M 31.5
RAP-SAM ConvNeXt-L [53] 52.3 82.5 58.6 42.8 61.1 55.6 700G 239M 8.6
Explore adapter Design on More Methods. We explore the effectiveness of our adapter design on Mask2Former [10] and add the adapter module to the last decoder layer of Mask2Former. As shown in Tab. 12, Mask2Former with adapter will improve 1.1 mIoU score compared with origin Mask2Former for interactive segmentation and only drop 0.3 PQ for panoptic segmentation. This result indicates that the adapter module not only applies to our method but can also be generalized to other methods
Table 12: Explore adapter Design on More Methods.
Method Backbone PQ mIoU Mask2Former R50 43.2 57.0 Mask2Former+adapter R50 43.0 58.1
More Implementation Details. We implement our proposed method and benchmark methods using MMDetection toolbox [5], and we train all of these methods in the same environment. For kMaX-DeepLab [85], we excluded the auxiliary semantic segmentation loss, instance discrimination loss, and preserving PQ-style loss alignment with Mask2former [10] for a fair comparison. We utilize the AdamW [54] optimizer with a weight decay of 0.05, and the learning rate is set to 1e-4 for all methods. We warm up the learning rate in the first 500 iterations using a linearly increased strategy and decay the learning rate at 8 and 11 epochs by a factor of 10. For data augmentation, we use the large-scale jittering (LSJ) augmentation [16] with a random scale sampled from the range of 0.1 to 2.0 and followed by a fixed size crop to 1024×1024.
More Inference Details. For video instance segmentation, we adopt simple object query matching [47, 26] for all methods to keep the fair comparison. For interactive segmentation, we mainly use point prompts, which is more challenging than box prompts. For panoptic segmentation, all hyperparameters are the same. The FPS is obtained on one A100 card for both the main paper and supplementary.
Appendix B Visual Comparison
Comparison on Youtube-VIS dataset. In Fig. 5, we compare our RAP-SAM (right) with a strong baseline Mask2Former (left) with the same ResNet50 backbone. Our methods achieve more consistent tracking and segmentation results in these examples.
Figure 5: When compared to the Mask2Former(right) on the YouTube-VIS 2019 dataset, our RAP-SAM(left) demonstrates superior performance in recognizing and segmenting objects in certain scenarios, and it also holds an advantage in detailing edges.
Comparison on COCO Panoptic Segmentation dataset. In Fig. 6, We demonstrate that our RAP-SAM model surpasses K-Net in panoptic segmentation tasks. RAP-SAM exhibits enhanced accuracy in processing complex scenes and maintaining fine details, especially in segmenting object edges and managing overlapping objects.
Figure 6: When compared with K-Net (bottom) on the COCO dataset, our RAP-SAM (top) model shows a significant advantage, achieving more accurate segmentation of foreground and background, with edges that are smoother.
More Interactive Segmentation Results. In Fig. 7, We present further visualizations of interactive segmentation, showcasing segmentation outcomes for ’things’ and ’stuff’. In these visualizations, green dots signify manually specified prompts. Our model exhibits precise segmentation abilities for ’things’ and ’stuff’. Notably, even when prompt points are situated on object boundaries, our model accurately identifies and segments the object.
Figure 7: Here are additional results of interactive segmentation. For point prompts or box prompts, RAP-SAM segments out the corresponding object based on the spatial location of the prompt. We have provided both segmentation results for ’things’ and ’stuff’, demonstrating that RAP-SAM is capable of segmenting both while maintaining strong performance.
Appendix C Challenges and Future Work
Failure Cases Analysis. We also analyze failure cases in video/image segmentation and identify two typical failure modes of RAP-SAM. Firstly, as illustrated in Fig. 8, when faced with multiple objects with high overlap, RAP-SAM struggles to fully recognize them, resulting in the omission of several masks. Secondly, in crowded scenarios, RAP-SAM faces challenges in recognizing and segmenting all instances with a limited number of instance kernels.
Figure 8: Failure modes of RAP-SAM in image/video segmentation.
Future Work Discussion. There are several directions to explore for real-time all-purpose segmentation. The first direction is to achieve better-balanced performance on image, video, and interactive segmentation (since there is still a lot of room for performance, as shown in the last row of Tab. 11). The second direction is to speed up the model and deploy the model on the edge device, such as iPhone. The third direction is to explore different knowledge distillation approaches to transfer the vision foundation model in real-time all-purpose models. Moreover, more visual prompts can be explored, such as mask prompts or box prompts. Our method currently only supports point prompts. This will be our future work.

### v2

@@
-RAP-SAM [Uncaptioned image]: Towards Real-Time All-Purpose Segment Anything
+RMP-SAM: Towards Real-Time Multi-Purpose Segment Anything
 Abstract
-Advanced by transformer architecture, vision foundation models (VFMs) achieve remarkable progress in performance and generalization ability.
-Segment Anything Model (SAM) is one remarkable model that can achieve generalized segmentation.
-However, most VFMs cannot run in real-time, which makes it difficult to transfer them into several products.
-On the other hand, current real-time segmentation mainly has one purpose, such as semantic segmentation on the driving scene.
-We argue that diverse outputs are needed for real applications.
-Thus, this work explores a new real-time segmentation setting, named all-purpose segmentation in real-time, to transfer VFMs in real-time deployment.
-It contains three different tasks, including interactive segmentation, panoptic segmentation, and video segmentation.
-We aim to use one model to achieve the above tasks in real-time.
-We first benchmark several strong baselines.
-Then, we present Real-Time All Purpose SAM (RAP-SAM).
-It contains an efficient encoder and an efficient decoupled decoder to perform prompt-driven decoding.
-Moreover, we further explore different training strategies and tuning methods to boost co-training performance further.
-Our code and model are available at https://github.com/xushilin1/RAP-SAM/.
+Recent segmentation methods, which adopt large-scale data training and transformer architecture, aim to create one foundation model that can perform multiple tasks.
+However, most of these methods rely on heavy encoder and decoder frameworks, hindering their performance in real-time scenarios.
+To explore real-time segmentation, recent advancements primarily focus on semantic segmentation within specific environments, such as autonomous driving.
+However, they often overlook the generalization ability of these models across diverse scenarios.
+Therefore, to fill this gap, this work explores a novel real-time segmentation setting called real-time multi-purpose segmentation.
+It contains three fundamental sub-tasks: interactive segmentation, panoptic segmentation, and video instance segmentation.
+Unlike previous methods, which use a specific design for each task, we aim to use only a single end-to-end model to accomplish all these tasks in real-time.
+To meet real-time requirements and balance multi-task learning, we present a novel dynamic convolution-based method, Real-Time Multi-Purpose SAM (RMP-SAM).
+It contains an efficient encoder and an efficient decoupled adapter to perform prompt-driven decoding.
+Moreover, we further explore different training strategies and one new adapter design to boost co-training performance further.
+We benchmark several strong baselines by extending existing works to support our multi-purpose segmentation.
+Extensive experiments demonstrate that RMP-SAM is effective and generalizes well on proposed benchmarks and other specific semantic tasks.
+Our implementation of RMP-SAM achieves the optimal balance between accuracy and speed for these tasks.
+The code is released at https://github.com/xushilin1/RAP-SAM
 1 Introduction
-Figure 1: We present real-time all-purpose segmentation to segment and recognize objects for image, video, and interactive inputs.
-In addition to benchmarking, we also propose a simple yet effective baseline, named RAP-SAM, which achieves the best accuracy and speed trade-off among three different tasks.
-Both real-time panoptic segmentation and video instance segmentation are shown at the bottom.
-Recent advancements in computer vision, propelled by transformer architectures [13, 2, 75], focus on developing a singular, large-scale model versatile enough for various tasks, including detection and segmentation, with a general-purpose design.
-Specifically, several studies [74, 80] adopting identical architectures have demonstrated superior performance compared to models designed exclusively for specific tasks.
-Furthermore, numerous studies [34, 30] investigate the foundation model design using extensive data.
-A notable example, the Segment Anything Model (SAM) [34], introduces an interactive model driven by visual prompts for versatile segmentation.
-Concurrently, for semantic-level segmentation, numerous studies [10, 93, 47, 46, 29] employing a universal design have surpassed previous models tailored for specific image and video segmentation tasks.
-However, the majority of these studies face challenges in achieving real-time performance or compatibility with mobile devices, primarily due to their heavy encoders and cascaded decoders.
-This constraint significantly limits the applicability of these advanced architectures in practical vision applications, although they can achieve high performance.
-On the contrary, various studies explore real-time segmentation using distinctive designs.
-Yet, these efforts predominantly concentrate on singular application purposes, including autonomous driving scenes, the COCO dataset for panoptic segmentation (PS) [50], and real-time video instance/object segmentation (VIS) [81].
-Currently, no studies investigate real-time, all-purpose segmentation, wherein a model is capable of performing universal segmentation tasks, encompassing image segmentation, video instance/video segmentation, and SAM-like interactive segmentation.
-Inspired by the aforementioned challenges, this study investigates the novel issue of real-time, all-purpose segmentation.
-We pose a critical inquiry: given the limited computational resources and model capacity, how can we develop an efficient, all-purpose segmentation model?
-This entails creating a single model capable of segmenting, tracking, and classifying each pixel in real-time, akin to performing interactive segmentation similar to SAM.
-Initially, we define this research problem by integrating existing benchmarks.
-Specifically, we use the COCO [50] and YouTube-VIS 2019 [81] datasets for joint co-training, employing identical hyper-parameters.
-Beyond the semantic queries used for semantic-level segmentation in prior research, our approach additionally incorporates visual prompt queries in SAM for interactive segmentation.
-However, there are still several challenges: 1), No previous works explore joint co-training for three different purposes.
-Thus, extensive experiments are needed to search for a better meta-architecture.
-2), Interactive segmentation and semantic-level segmentation need better strategies to balance each other.
-3), Under the real-time setting, several previous architectures [80, 42] for multi-task segmentation may not be practical.
-To address these challenges, we benchmark several recent works [85, 10, 25] in a real-time setting.
-Then, we initially investigate a range of meta-architectures suitable for all-purpose segmentation.
-Unlike previous works, we avoid cascaded architecture and only use the pyramid feature once when performing dynamic convolution.
-We explore four different architectures in the case of various decoder designs, where we adopt a single decoder with multiple queries.
-Via extensive experiments, we adopt pooling-based dynamic convolution to replace per-pixel cross-attention to achieve better accuracy and speed trade-off.
-To balance interactive and semantic-level segmentation, we present an effective dual adapter design that better adapts shared knowledge of the same decoder.
-Finally, we term our best model as Real-time All-Purpose SAM (RAP-SAM).
-To the best of our knowledge, it is the first real-time all-purpose segmentation model.
-Moreover, due to the semantic level and interactive level segmentation ability, our model can also create more new applications, such as interactive video segmentation.
+Recent advancements in computer vision, driven by transformer architectures (Dosovitskiy et al., 2021; Carion et al., 2020; Wang et al., 2023b), focus on creating a single large-scale model versatile enough to handle different tasks like detection and segmentation with a general-purpose design.
+Several studies utilizing the same architectures (Wang et al., 2023a; Yan et al., 2023; Jain et al., 2023) have shown superior performance across various tasks.
+Enjoying the benefit of the large-scale model capacity and training data, the general-purpose models demonstrate strong performance on versatility for specific applications (Kirillov et al., 2023; Jia et al., 2021).
+For image and video segmentation at the semantic level, numerous studies (Cheng et al., 2022; Zhang et al., 2021; Li et al., 2022d; 2023e; Jain et al., 2023) employing a universal design have surpassed previous models tailored for specific image and video segmentation tasks.
+The Segment Anything Model (SAM) (Kirillov et al., 2023) is a leading approach in interactive segmentation, offering a unified model for various interactions such as point, bounding box, mask, and text input.
+Figure 1: We present real-time multi-purpose segmentation to segment and recognize objects for image, video, and visual prompt inputs.
+In addition to benchmarking, we propose a simple yet effective baseline, named RMP-SAM, which achieves the best performance and speed trade-off among three different tasks.
+The larger dot indicates more parameters.
+However, most of these studies face challenges in achieving real-time performance on devices with limited computing resources due to their heavy encoders or cascaded decoders.
+Despite their high-performance capabilities and versatility, this limitation hinders the practical use of these advanced architectures in vision applications, such as editing tools on edge devices.
+In contrast, several studies (Zhao et al., 2018; Hu et al., 2023) focus on real-time segmentation using specific designs (e.g., efficient decoder fusion) but are often tailored to a specific application trained on a particular dataset (e.g., COCO) or restricted to similar tasks like video semantic segmentation.
+To our knowledge, no study investigates real-time, multi-purpose segmentation, where a single model can perform universal segmentation tasks, including image, video, and SAM-like interactive segmentation.
+For example, given a video input, we want to create a tool to perform various segmentation tasks on the device, including interactive segmentation with visual cues, labeling the masks, and tracking the masks.
+Inspired by the abovementioned challenges, we focus on real-time multi-purpose segmentation and efficiently investigate the novel issue of training versatile models.
+We pose a critical inquiry: Given the limited computational resources and model capacity, how can we develop an efficient, multi-purpose segmentation model?
+This entails creating a single model capable of segmenting, tracking, and classifying each pixel in real-time, similar to performing interactive segmentation similar to SAM.
+We present an illustration figure in Fig.
+1(a).
+First, we explore this problem by benchmarking existing methods.
+Specifically, we use the COCO (Lin et al., 2014) and YouTube-VIS (Yang et al., 2019) for joint co-training, employing identical hyperparameters.
+Beyond the object queries used for image and video segmentation in prior research, our approach incorporates visual prompt queries for interactive segmentation.
+However, none of these works can achieve the best accuracy and speed trade-off on three tasks using one model.
+This is because: 1) No previous works explore joint co-training for segmentation for different purposes.
+Thus, extensive experiments are needed to gain a better understanding of meta-architecture.
+2) Interactive and semantic-level segmentation needs better strategies to balance each other.
+3) Under the real-time setting, several previous architectures (Yan et al., 2023; Li et al., 2022c) for multi-task learning may not be practical.
+To solve these issues, we present a real-time multi-purpose segment anything model (RMP-SAM).
+It contains an efficient encoder and shared decoder with multiple queries as inputs.
+Unlike previous works, we avoid cascaded transformers encoder layers and pixel-wise cross-attention decoder (Cheng et al., 2022).
+Instead, we propose a pooling-based dynamic convolution to replace per-pixel cross-attention to achieve better accuracy and speed trade-off.
+We represent all three tasks using queries with shared dynamic convolution operation.
+According to different outputs, we can divide the segmentation tasks into two types: one for class-agnostic masks (SAM-like) and the other for class-specific masks (contains one mask and one class label).
+Two types have different goals and will conflict during the multi-task learning process.
+Thus, we further explore four different architectures in the case of various decoder designs.
+In particular, we find a single decoder with multiple queries for different tasks can achieve good enough results and best speed and accuracy trade-off.
+Moreover, we present an effective dual adapter design that better adapts shared knowledge of the same decoder.
+The dual adapter contains an asymmetric design to better transfer shared features.
+To the best of our knowledge, RMP-SAM is the first real-time multi-purpose segmentation model that can achieve the best speed and accuracy trade-off on four different tasks, as shown in Fig.
+1(b).
+Moreover, due to the semantic level and interactive level segmentation ability, our model can create more new applications, such as interactive video segmentation.
+In addition, we also verify our model in the SAM-like setting by co-training with SAM-data, it also achieves comparable results.
 Extensive experiments show the effectiveness of our design.
 The main contributions of this work are:
-  • We introduce all-purpose segmentation, a multi-task segmentation that aims to segment objects for image, video, and interactive inputs in real-time.
+  • We introduce real-time multi-purpose segmentation, a novel multi-task segmentation aiming to segment objects for real-time image, video, and interactive inputs.
   • We benchmark several real-time transformer-based segmentation approaches for the new settings.
-  • We present a simple yet fast baseline named RAP-SAM. It contains a lightweight feature extractor, a unified decoder, and two asymmetric adapters.
-  • Extensive experiments demonstrate that RAP-SAM achieves the best speed and accuracy trade-off in the proposed benchmark and regular real-time semantic and panoptic segmentation benchmarks.
+  • We present a simple yet fast baseline named RMP-SAM. It contains a lightweight feature extractor, a unified decoder, and two asymmetric adapters.
+  • Extensive experiments demonstrate that RMP-SAM achieves the best speed and accuracy trade-off in the proposed benchmark and regular real-time semantic and panoptic segmentation benchmarks.
 We also show scalability across datasets and application demos.
-2 Related Work
-Universal Segmentation.
-Prior research has focused on designing segmentation models tailored for specific tasks.
-Recent developments in this field [13, 52, 2, 93, 10, 85, 49, 39, 17, 86, 42, 43, 45, 87] have shifted towards employing mask classification architectures, coupled with an end-to-end set prediction objective, to facilitate universal segmentation.
-This approach has outperformed specialized models [4, 33, 20, 48, 27, 88, 96, 7, 40, 41, 6] across various segmentation tasks, including those involving images, videos, and point clouds [32, 47, 46, 9, 67, 78, 79].
-Specifically, Mask2Former [10] uses a masked-attention mechanism and surpasses the performance of prior specific image segmentation models.
-Similarly, Tube-Link [46] outperforms earlier models [47, 40, 51] tailored for specific video segmentation tasks.
-Semantic-SAM [35], built on Mask-DINO [36], employs a hybrid query design to facilitate both panoptic and interactive segmentation within a single framework.
-However, despite the strong performance of these works, their complexity, stemming from the heavier encoder and the cascaded decoder designs, poses significant challenges for integration into actual products.
-These are primarily hindered by slow processing speeds and substantial parameter overhead.
-Vision Foundation Model.
-Recent advancements in the vision research community have been propelled by the development of Vision Foundation Models (VFMs), including pure-visual pre-training [19], vision-language pre-training [15, 37], and multi-modal models employing visual prompting [75, 1].
-SAM [34] stands as a seminal work targeting general interactive segmentation.
-Subsequent studies [3, 12, 77] have adapted SAM for diverse applications, predominantly employing it as a segmentation tool.
-However, these Vision Foundation Models typically employ substantial visual encoders, like ViT-base [13], which poses challenges for deployment on real-time or mobile devices due to their size.
-Efficient Model Design.
-This research direction primarily concentrates on the development of efficient CNNs [24, 66, 23, 28, 94, 56, 18], transformers [57, 64, 58], and hybrid architectures [90, 55, 8, 38, 100], aimed at advancing visual representation learning.
-Most research investigates backbone design within the constraints of real-time processing and parameter efficiency.
-These studies are orthogonal to our own, which primarily explores the impact of data, task, and decoder design.
-Additionally, our work employs efficient models as encoders and presents detailed benchmarks and analyses of the encoder’s effects.
-Efficient Segmentation.
-Previous studies [95, 44, 84, 25, 22, 71, 83, 59] on efficient segmentation have predominantly concentrated on closed-set and specific domains.
-In particular, a significant portion of this research [44, 25, 60] is dedicated to driving scenarios.
-Recently, various studies [92, 71, 99] have developed efficient segmentation techniques that facilitate model execution on mobile devices.
-Mobile SAM [89] introduces a streamlined encoder distillation method.
-Fast SAM [97] employs a single-stage instance segmentation framework that directly decodes class-agnostic masks.
-Recently, multiple studies have been conducted on efficient panoptic segmentation [69, 25, 63] and rapid video instance segmentation [82, 91].
-However, these real-time segmentation methods are limited to specific tasks.
-We contend that an all-purpose model capable of real-time performance would have wide-ranging applications in editing, tracking, and segmentation functionalities within various products.
-3 Problem Setup
-Settings.
-Our proposed real-time all-purpose segmentation contains three visual inputs, including image, video, and visual prompts (boxes and points).
-It outputs corresponding masks, labels, and instance ID.
-We adopt panoptic segmentation for image segmentation, which unifies semantic and instance segmentation.
-For video segmentation, we adopt video instance segmentation, where we aim to segment and track each foreground instance.
-For interactive segmentation, we follow the SAM-like setting, where we perform class-agnostic segmentation according to the user’s inputs.
-Compared with previous settings, as shown in Tab.
-1, we can achieve more flexible input and output, which is more challenging.
-Combined with interactive segmentation and video segmentation, we can also achieve interactive video segmentation as video object segmentation.
-Datasets.
-For image segmentation and interactive segmentation, we use the well-known COCO dataset [50] for benchmarking.
-We adopt partially SAM data [34] for testing.
-Due to the limited resources and limited model capacity in real-time settings, we do not use SAM data for training.
-We take COCO data as SAM data for training and testing, where the boxes and random points sampled from masks are used for visual prompting.
-For video segmentation, we adopt the widely used YouTube-VIS 2019 dataset [81] for training.
-We do not use the SAM data for our benchmarking.
-There are several reasons: (1), SAM data is not well annotated as the COCO dataset.
-(2), We only focus on object-level and scene-level segmentation, while SAM data contain more granularity.
-(3), SAM data has no semantics, while our goal is to predict mask labels for real applications.
-Figure 2: RAP SAM overview.
-Our method contains three visual inputs: image, video, and visual prompts.
-Utilizing positional encoding, we generate prompt queries from these visual prompts.
-The learnable object queries, alongside the prompt queries and the feature map $F$, are directed to the multi-stage decoder.
-This process generates multi-stage predictions and refined queries.
-These refined queries engage in cross-attention with $F$, resulting in the final prediction.
-4 Proposed Method
-Motivation.
-Employing the same formulation, we first revisit three tasks: panoptic segmentation, video instance segmentation, and interactive segmentation.
-When combined with a query-based mask transformer, object queries can represent all entities within the scene, encompassing image-level objects, video-level tracked objects, and prompt-based specific objects.
-Each task’s formulation is as follows:
-For panoptic segmentation [33], given an input image $I\in\mathbb{R}^{H\times{W}\times 3}$, the objective is to generate a group of masks $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i})\}_{i=1}^{N}\,$ where $c_{i}$ denotes the ground truth class label of the binary mask $m_{i}$ and $N$ is the number of masks, ${H}\times{W}$ indicates the spatial dimensions.
-For video instance segmentation [81], given a video clip input $V\in\mathbb{R}^{T\times H\times{W}\times 3}$, where $T$ denotes the number of frames, the process aims to generate an instance mask tube $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i},d_{i})\}_{i=1}^{N}\,$, where $N$ represents the number of the tube masks $m_{i}\in{\{0,1\}}^{{T}\times{H}\times{W}}$, $c_{i}$ indicates the class label of the tube $m_{i}$, and $d_{i}$ identifies the instance ID for each tube.
-Table 1: Real-Time Segmentation Settings. Our all-purpose segmentation supports more tasks in one framework.
-Property Image Seg Video Seg SAM-Like Ours
+2 Real-Time Multi-Purpose Segmentation
+Table 1: Real-Time Segmentation Settings. Our multi-purpose segmentation supports more tasks in one framework.
+Property Image Seg Video Seg SAM-Like SAM-2-Like Ours: Real-Time Multi-Purpose Seg
 Image Masks
@@
 Multitasks
-Table 2: Comparison of Segmentation Methods. Our proposed RAP-SAM supports various segmentation tasks, and it can run in real time.
+Table 2: Comparison of Segmentation Methods. Our proposed RMP-SAM supports various segmentation tasks and performs them in real-time.
 Methods SS PS VIS Interactive Multi-Task in One Model Real Time
-ICNet [95]
-Bi-Seg [84]
-YOSO [25]
-Mobilie-VIS [91]
-SAM [34]
-Mask2Former [10]
-Video K-Net [47]
-OneFormer [29]
-RAP-SAM (Ours)
-For SAM-like interactive segmentation [34], both the image $I$ and visual prompts $P\in\mathbb{R}^{K\times 4}$ (box prompts are used here) are taken as inputs.
+ICNet (Zhao et al., 2018)
+Bi-Seg (Yu et al., 2018)
+YOSO (Hu et al., 2023)
+Mobilie-VIS (Zhang et al., 2024)
+SAM (Kirillov et al., 2023)
+Mask2Former (Cheng et al., 2022)
+Video K-Net (Li et al., 2022d)
+OneFormer (Jain et al., 2023)
+RMP-SAM (Ours)
+Problem Settings.
+Our proposed real-time multi-purpose segmentation contains three visual inputs: image, video, and visual prompts (boxes and points).
+It predicts the corresponding masks, labels, and instance IDs for the content within the given input.
+Basically, we adopt the definition of panoptic segmentation Kirillov et al.
+(2019) for image inputs since it covers all concepts in the scene.
+For video inputs, we segment and track each foreground instance (video instance segmentation).
+Given visual prompts, we follow the SAM-like Zhang et al.
+(2023a); Zhou et al.
+(2023); Shu et al.
+(2023); Xiong et al.
+(2023) setting for interactive segmentation, performing class-agnostic segmentation according to the user’s inputs.
+Although recent advancements, such as SAM-2 (Ravi et al., 2024), have attempted to unify the interactive segmentation task for both image and video inputs, our framework emphasizes the importance of object semantics.
+It incorporates multi-task learning in a single shot.
+Combined with interactive segmentation and video segmentation, we can also achieve interactive video segmentation as video object segmentation.
+As shown in Tab.
+1 and Tab.
+2, we compare previous segmentation methods with and indicate that our method unifies different segmentation tasks within a single framework.
+This approach enables more flexible usage across various segmentation tasks, particularly regarding input modalities and predictions.
+Datasets.
+For benchmarking, we use the well-known COCO dataset (Lin et al., 2014) for panoptic and interactive segmentation.
+During the training of interactive segmentation, the boxes derived from the ground truth, along with points randomly sampled from the masks, serve as visual prompts.
+For testing, we obtain the center point of the mask as a visual prompt.
+For video segmentation, we adopt the widely used YouTube-VIS 2019 dataset (Yang et al., 2019) for training.
+We do not use the SAM data for our benchmarking.
+There are several reasons: (1) SAM data is not as well annotated as the COCO dataset.
+(2) We only focus on object-level and scene-level segmentation, while SAM data contain more granularity.
+(3) SAM data has no semantics, while our goal is to predict semantic labels for real applications.
+To fairly compare with SAM-like models, we adopt the SAM-data training for specific setting, as shown in Sec.
+4.1.
+3 Proposed Method
+First, we revise panoptic segmentation, video instance segmentation, and interactive segmentation under the unified formulation.
+When combined with a query-based segmentation model, object queries can represent all entities within the scene, encompassing image-level objects, video-level tracked objects, and prompt-based specific objects.
+For panoptic segmentation (Kirillov et al., 2019), given an input image $I\in\mathbb{R}^{H\times{W}\times 3}$, the objective is to generate a group of masks $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i})\}_{i=1}^{N}\,$ where $c_{i}$ denotes the class label of the binary mask $m_{i}$ and $N$ is the number of queries, ${H}\times{W}$ indicates the spatial dimensions.
+For video instance segmentation (Yang et al., 2019), given a video clip input $V\in\mathbb{R}^{T\times H\times{W}\times 3}$, where $T$ denotes the number of frames, the process aims to generate an instance mask $\{y_{i}\}_{i=1}^{N}=\{(m_{i},c_{i},d_{i})\}_{i=1}^{N}\,$, where $N$ represents the number of the tube masks $m_{i}\in{\{0,1\}}^{{T}\times{H}\times{W}}$, $c_{i}$ indicates the class label of the tube $m_{i}$, and $d_{i}$ identifies the instance ID for each instance.
+For SAM-like segmentation (Kirillov et al., 2023), both the image $I$ and visual prompts $P\in\mathbb{R}^{K\times 2}$ (point prompts are used here) are taken as inputs.
 These prompts, which include points and boxes, lead to the outputs of corresponding binary image masks $\{y_{i}\}_{i=1}^{K}=\{m_{i}\in\{0,1\}^{H\times W}\}_{i=1}^{K}$, where $K$ denotes the number of visual prompts.
-Each visual prompt is encoded into one object query, which naturally can be the input of the decoder, as implemented in  [34, 10].
-Single Object Query For All Segmentation.
-As mentioned above, by combining all three settings into one, we can represent all the output segmentation entities using the same query-based mask classification framework.
-In particular, one object query corresponds to one mask $m_{i}$, one label $c_{i}$, and one ID $d_{i}$.
-We aim to design a simple, efficient, and parameter-sharing framework to support these tasks.
-Compared Baselines.
-In Sec.
-5.1, we benchmark several baselines for our all-purpose segmentation.
-We use the prompt encoder (proposed in SAM [34]) to encode the visual prompt and jointly learn with object query for both PS and VIS on a shared decoder.
-Although we adopt the lightweight design to speed up, their performance is still limited or not balanced, as shown in Tab 3.
-4.1 RAP-SAM Architecture
+Each visual prompt is encoded into one object query, which naturally can be the input of the decoder.
+3.1 RMP-SAM Architecture
 Overall Architecture.
-Our RAP-SAM is a simple encoder and decoder architecture.
 As shown in Fig.
-2, it contains a backbone, a lightweight neck, and a shared multitask decoder.
-Visual prompts $P$ are also the input of the decoder.
-Following SAM, we also adopt the prompt encoder to encode visual prompts into a query.
-We adopt the same decoder for both visual prompts and initial object queries to share more computation and parameters.
-However, the goals of both are different.
-The former pays more attention to the local details, while the latter also considers the scene and temporal features.
-To better balance the results for interactive segmentation and image/video segmentation, we design a prompt adapter and an object adapter in the end of the decoder.
+2, our RMP-SAM is built upon a simple encoder and decoder architecture that can input images, videos, and visual prompts.
+Following SAM, we utilize the prompt encoder to convert visual prompts into queries.
+These visual prompts and object queries share the same decoder for the efficiency of computation and parameter size.
+However, the goals of the two types of queries are different: the former pays more attention to the local details of the visual prompts, while the latter considers the scene and temporal features.
+To better balance the results for interactive segmentation and image/video segmentation, we design a prompt adapter and an object adapter at the end of the decoder.
+Figure 2: RMP-SAM overview.
+Our method contains three visual inputs: image, video, and visual prompts.
+Utilizing positional encoding, we generate prompt queries from these visual prompts.
+The learnable object queries, prompt queries, and the feature map $F$ are directed to the multi-stage decoder.
+This process generates multi-stage predictions and refined queries.
+These refined queries engage in cross-attention with $F$, resulting in the final prediction.
 Lite Feature Extractor.
-Due to computation cost limitation, we avoid the design of previous methods, such as a large backbone and heavier transformer encoder.
-We explore lightweight backbones, including ResNet18 [21], STDC-v1 [14], and SeaFormer [71].
-We adopt feature pyramid networks with deformable convolutions to fuse multi-scale features and obtain more aligned feature representation, motivated by previous works [44] on real-time semantic segmentation.
-For video input, we extract the spatial-temporal features.
+Due to computation cost limitations, we avoid using previous methods (Cheng et al., 2022; Li et al., 2024b) that involve large backbones and heavier encoders.
+As required by real-time constraint, we investigate lightweight backbones such as ResNet18 (He et al., 2016) and TopFormer (Zhang et al., 2022).
+We incorporate a feature pyramid network with deformable convolution to enhance representation alignment and fuse multi-scale features in real-time.
+When dealing with video input, we extract the spatial-temporal features.
 For simplicity, we use $F_{img}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times d}$ for image inputs and $F_{vid}\in\mathbb{R}^{T\times\frac{H}{4}\times\frac{W}{4}\times d}$ for video inputs.
 Unified Dynamic Convolution Decoder.
-As in previous works, the goal of the decoder is to refine the object query.
-However, many of these approaches are impractical for real-time settings due to their reliance on heavily cascaded layers and pixel-wise cross-attention mechanisms.
-In contrast, our method employs a pooling-based dynamic convolution framework [93, 47] to enhance the decoder’s efficiency.
-Given object query $Q_{i}\in\mathbb{R}^{N\times d}$ and prompt query $P_{i}\in\mathbb{R}^{K\times d}$, the initial masks $M_{i}$ are obtained by dot product with $F_{img}$ or $F_{vid}$.
-$i$ denotes the index of the decoder stage.
+The goal of the decoder is to refine the object query.
+However, many of these methods (Cheng et al., 2021b; 2022), are not feasible for real-time settings due to their reliance on heavily cascaded layers and pixel-wise cross-attention mechanisms.
+In contrast, our method uses a pooling-based dynamic convolution framework (Zhang et al., 2021; Li et al., 2022d) to enhance the efficiency of the decoder.
+Given object query $Q\in\mathbb{R}^{N\times d}$ and prompt query $P\in\mathbb{R}^{K\times d}$, the masks $M_{i}$ are obtained by dot product with $F_{img}$ or $F_{vid}$.
+Here, $i$ represents the decoder stage index, $N$ is the total number of object queries, and $K$ is the number of visual prompts.
 According to the inputs, the masks can be image masks for panoptic segmentation $M_{i}^{pan}\in\mathbb{R}^{N\times\frac{H}{4}\times\frac{W}{4}}$ and interactive segmentation $M_{i}^{iter}\in\mathbb{R}^{K\times\frac{H}{4}\times\frac{W}{4}}$ or tube masks $M_{i}^{tube}\in\mathbb{R}^{N\times T\times\frac{H}{4}\times\frac{W}{4}}$.
@@
 In our implementation, the learned query is shared for both $Q_{i}^{pan}$ and $Q_{i}^{tube}$.
-In particular, we adopt the design [93, 68, 42] to refine input queries $Q_{i}^{pan}$ with features $X_{i}^{pan}$ which are grouped from their masks,
+In particular, we adopt the design of previous methods (Zhang et al., 2021; Sun et al., 2021; Li et al., 2022c) to refine input queries $Q_{i}^{pan}$ with features $X_{i}^{pan}$ which are grouped from their masks,
 $\hat{Q}^{pan}_{i}=\mathrm{DynamicConv}(X_{i}^{pan},Q^{pan}_{i-1}),$ (3)
@@
 Specifically, $\mathrm{DynamicConv}$ uses input query features $X^{pan}_{i}$ to generate gating parameters via a multilayer perceptron (MLP) and multiply back to the original query input $Q^{pan}_{i-1}$.
-We adopt the same design [93, 68] by learning gating functions to update the refined queries.
+We adopt the same design (Zhang et al., 2021; Sun et al., 2021) by learning gating functions to update the refined queries.
 The $\mathrm{DynamicConv}$ operation is based on:
 $\hat{Q}^{pan}_{i}=\mathrm{Gate}_{x}(X^{pan}_{i})X^{pan}_{i}+\mathrm{Gate}_{q}(X^{pan}_{i})Q^{pan}_{i-1},$ (4)
-where $\mathrm{Gate}$ is implemented with a fully connected (FC) layer followed by Layer Norm (LN) and a sigmoid layer.
+where $\mathrm{Gate}$ is implemented with a fully connected layer followed by Layer Norm and a sigmoid layer.
 We adopt two different gate functions, including $\mathrm{Gate}_{x}$ and $\mathrm{Gate}_{q}$.
-They are implemented by MLP.
 The former is to weigh the query features, while the latter is to weigh corresponding queries.
-Next, we adopt one self-attention layer with feed-forward layers [70, 72] to learn the correspondence among each query and update them accordingly.
+Next, we adopt one self-attention layer with feed-forward layers (Vaswani et al., 2017; Wang et al., 2021) to learn the correspondence among each query and update them accordingly.
 This operation leads to the full correlation among queries, shown as follows:
 $Q_{i}=\mathrm{FFN}(\mathrm{MHSA}(\hat{Q}_{i})+\hat{Q}_{i}),$ (5)
-where $\mathrm{MHSA}$ means Multi Head Self Attention, $\mathrm{FFN}$ is the Feed Forward Network commonly used in the current version of Transformers [2, 13].
-Finally, the refined masks are obtained via dot product between the refined queries $Q^{pan}_{i}$, and the input features $F_{p}$, $F_{s}$.
-For mask classification, we adopt several feed-forward layers on $Q^{pan}_{i}$ and directly output the class scores.
-Then, we perform the inner product between the learned queries $Q^{pan}_{i}$, $Q^{iter}_{i}$, and their features $F_{vid}$ and $F_{img}$.
-For the video segmentation task, the only difference is that we predict the tube mask rather than the image mask.
+where $\mathrm{MHSA}$ means Multi Head Self Attention, $\mathrm{FFN}$ is the Feed Forward Network commonly used in the current version of Transformers (Carion et al., 2020; Dosovitskiy et al., 2021).
+We adopt several feed-forward layers on $Q^{pan}_{i}$ for mask classification and directly output the class scores.
+Then, we perform the inner product between the learned queries $Q^{pan}_{i}$, and image features $F_{vid}$ and $F_{img}$.
+The only difference for the video segmentation task is that we predict the tube mask rather than only a single image mask.
 The entire process (Equ.
@@
 4.2 for the detailed decoder design.
-Table 3: Real-Time All-Purpose Segmentation benchmark. Our proposed RAP-SAM achieves the best accuracy and speed trade-off on three tasks.
-Method Backbone COCO-Panoptic COCO-SAM YouTube-VIS 2019 Flops Parameters FPS
-PQ SQ $\mathrm{PQ_{th}}$ $\mathrm{PQ_{st}}$ mIoU mAP
-Mask2Former [10] R18 35.6 77.5 39.0 30.3 54.7 38.6 89.8G 18.6M 31.2
-Mask2Former [10] R50 42.9 79.8 47.6 35.6 58.0 42.1 153G 45.2M 26.6
-MaskFormer [11] R18 31.0 76.0 33.2 27.7 55.6 34.1 79.7G 18.5M 38.0
-MaskFormer [11] R50 37.4 77.3 41.2 31.8 58.8 40.0 143.0G 45.2M 34.3
-MaskFormer [11] TopFormer [92] 31.6 75.9 33.5 28.6 56.1 38.2 60.0G 12.8M 29.9
-kMaX-DeepLab [85] R18 27.8 71.0 30.3 24.1 16.9 23.1 87.1G 18.7M 15.0
-kMaX-DeepLab [85] R50 36.9 75.7 42.7 28.4 20.1 26.4 280.0G 51.4M 14.9
-YOSO [25] R18 31.6 76.8 36.4 24.4 45.7 31.4 57.3G 18.7M 41.0
-YOSO [25] R50 37.3 76.9 42.7 29.2 49.2 40.7 119.0G 45.1M 36.3
-YOSO [25] TopFormer [92] 31.0 75.9 33.0 26.9 45.1 27.9 36.8G 12.9M 31.1
-RAP-SAM R18 39.9 78.6 43.3 34.8 52.7 38.7 60.5G 22.8M 40.3
-RAP-SAM R50 46.9 80.8 51.6 39.8 57.9 46.2 123.0G 47.2M 35.1
-RAP-SAM TopFormer [92] 34.6 77.0 37.8 29.8 53.3 41.7 40.1G 15.0M 30.7
 Light-Weight Decoupled Adapter.
@@
 $A_{obj}$ uses the same dynamic convolution design to refine the object query further, while $A_{prompt}$ uses pixel-wise cross-attention design.
-Our key insights are: (1) For panoptic segmentation and video segmentation, the interaction of the scene feature is necessary, but not for interactive segmentation.
+Our key insights are: (1) The interaction between the whole scene feature and object queries is necessary for panoptic and video segmentation but not for interactive segmentation.
+The visual prompts provide strong positional information, and they pay more attention to local details features.
 (2) Despite the decoder being shared, the details are still missing for interactive segmentation due to the pooling effect in Equ.
@@
 More detailed adapter design can be found in Sec.
-5.2.
-4.2 Training and Inference
+4.2.
+3.2 Training, Inference, and Extension
 Joint Image and Video Segmentation Co-training.
 Our goal is to train all segmentation tasks only once jointly.
-All training targets are one entity label and mask for all three different cases.
-The entity can be thing, stuff, class-agnostic masks, and their corresponding labels.
+All training targets are one entity label and mask for all three cases.
+The entity can be thing, stuff, class-agnostic masks, and corresponding labels.
 Note that the instance masks with the same ID $d$ form the tube masks.
 During training, we apply Hungarian matching between the predicted and ground-truth entity masks to assign object queries to video/image entities and then supervise their predicted masks and classification.
-The classifier is replaced by CLIP text embedding to avoid cross-dataset taxonomy conflicts.
-Following previous works [10, 46], the final loss function is given as $L=\lambda_{cls}L_{m\_cls}+\lambda_{ce}L_{m\_ce}+\lambda_{dice}L_{m\_dice}$.
-Here, $L_{m\_cls}$ is the Cross-Entropy (CE) loss for mask classification, and $L_{m\_ce}$ and $L_{m\_dice}$ are mask Cross-Entropy (CE) loss and Dice loss [62, 73] for segmentation, respectively.
+To unify label taxonomies across datasets, we replace the learnable classifier with CLIP text embeddings following ViLD (Gu et al., 2021).
+Following previous works (Cheng et al., 2022; Li et al., 2023e), the final loss function is given as $L=\lambda_{cls}L_{m\_cls}+\lambda_{ce}L_{m\_ce}+\lambda_{dice}L_{m\_dice}$.
+Here, $L_{m\_cls}$ is the Cross-Entropy (CE) loss for mask classification, and $L_{m\_ce}$ and $L_{m\_dice}$ are mask Cross-Entropy (CE) loss and Dice loss (Milletari et al., 2016; Wang et al., 2020) for segmentation, respectively.
 By default, we set $\lambda_{cls}=2,\lambda_{ce}=5,\lambda_{dice}=5$.
 Inference.
-we follow the same inference procedure of Mask2Former [10] for image segmentation.
-We merge the things and stuff for PS according to the sorted scores.
-The scores are generated by CLIP text embedding.
-For video instance segmentation, we use query matching rather than introducing extra tracking components to generate instance ID, following previous work [47, 26, 46].
+We follow the same inference procedure of Mask2Former (Cheng et al., 2022) for image segmentation.
+In panoptic segmentation, we merge the prediction of things and stuff according to the sorted scores.
+For video instance segmentation, we use query matching rather than introducing extra tracking components to generate instance ID, following previous work (Li et al., 2022d; Huang et al., 2022; Li et al., 2023e).
 We adopt a near-online method for inference.
-For interactive segmentation tasks, we follow the origin SAM [34], by providing box and point prompts and obtaining the binary masks.
+We follow the original SAM (Kirillov et al., 2023) for interactive segmentation tasks by providing box and point prompts and obtaining the binary masks.
 All the parameters are shared across three different tasks.
 Extension and Application.
-Since our model can perform various segmentation tasks, RAP-SAM can be extended to more datasets and tasks, including semantic segmentation on ADE20k [98] or video panoptic segmentation on VIPSeg [61].
+Since our model can perform various segmentation tasks, Our method can be extended to more datasets and tasks, including semantic segmentation on ADE20k (Zhou et al., 2017) or video panoptic segmentation on VIPSeg (Miao et al., 2022).
 Moreover, our method can support prompt-driven video object segmentation by combining interactive segmentation with a video instance segmentation framework.
-We put the details of the extension and application for RAP-SAM in the supplementary.
+We put the details of the extension and application for RMP-SAM in the appendix (Sec.
+7.1).
 4 Experiments
 Datasets.
-Our benchmark uses COCO [50] and YouTube-VIS 2019 [81] datasets for benchmarking.
-COCO-SAM is constructed by using ground truth boxes and masks as visual prompt inputs, following the SAM-1B dataset [34].
-In addition, to verify the effectiveness and generality of RAP-SAM, we also use other datasets, including ADE-20K [98] and VIP-Seg dataset [61] in Sec.
-5.1.
+Our benchmark uses COCO (Lin et al., 2014) and YouTube-VIS 2019 (Yang et al., 2019) datasets for benchmarking.
+COCO-SAM is constructed using ground truth boxes and points randomly sampled from ground truth masks as visual prompt inputs.
+In addition, to verify the effectiveness and generality of RMP-SAM, we also use other datasets, including ADE-20K (Zhou et al., 2017) and VIP-Seg dataset (Miao et al., 2022) in Sec.
+4.1.
 Evaluation Metrics and Devices.
 For image segmentation, we adopt panoptic quality (PQ), which can be further decomposed into the segmentation quality (SQ) and the recognition quality (RQ).
-For the VIS task, mAP is adopted.
-We further report the semantic segmentation results (mIoU) and video panoptic segmentation results (STQ [76] and VPQ [31]).
-We adopt one A100 GPU for speed testing.
+For the VIS task, we use mAP as our primary metric.
+We further report the semantic segmentation results (mIoU) and video panoptic segmentation results (STQ (Weber et al., 2021) and VPQ (Kim et al., 2020)).
+Speed testing is conducted fairly on one A100 GPU for all models.
 Re-implemented Baselines.
-Since no previous works explore joint co-training for image, video, and interactive segmentation in one framework.
-We re-implement several representative baselines using the same codebase, including non-real-time models (Mask2Former [10], kMaX-DeepLab [85]) for reference.
-In particular, we benchmark over ten different methods, including different backbones and decoders.
+Since no previous works explore joint co-training for image, video, and interactive segmentation in one model.
+We re-implement several representative baselines using the same codebase, including non-real-time models (Mask2Former (Cheng et al., 2022), kMaX-DeepLab (Yu et al., 2022)) for reference.
+We benchmark multiple different methods, including different backbones and decoders.
 Implementation Details.
-We implement our models and all other baselines in PyTorch [65] with the MMDetection toolbox [5].
+We implement our models and all other baselines in PyTorch (Paszke et al., 2019).
 We use the distributed training framework with 16 A100 GPUs.
-Each mini-batch has two images per GPU.
+Each mini-batch has two images per GPU, and each batch contains one data type.
 In particular, we adopt pseudo video training on COCO by moving image masks with random directions.
-All the models are trained with 24 epochs.
-For data augmentation, we adopt large-scale jitter as previous works [10] to build strong baselines.
-For all models, we adopt training steps and optimizers.
-Refer to the supplementary for more details.
+All the models are trained with 12 epochs.
+For data augmentation, we adopt large-scale jitter as previous works (Cheng et al., 2022) to build strong baselines.
+For all models, we adopt the same training steps and optimizers.
+Refer to the appendix (Sec.
+7.1) for more details.
+Table 3: Real-Time Multi-Purpose Segmentation benchmark. Our proposed RMP-SAM achieves the best accuracy and speed trade-off on three tasks.
+Method Backbone COCO-Panoptic COCO-SAM YouTube-VIS 2019 FLOPs Parameters FPS
+PQ SQ $\mathrm{PQ_{th}}$ $\mathrm{PQ_{st}}$ mIoU mAP
+Mask2Former R18 35.6 77.5 39.0 30.3 54.7 38.6 89.8G 18.6M 31.2
+MaskFormer R18 31.0 76.0 33.2 27.7 55.6 34.1 79.7G 18.5M 38.0
+kMaX-DeepLab R18 27.8 71.0 30.3 24.1 16.9 23.1 87.1G 18.7M 15.0
+YOSO R18 31.6 76.8 36.4 24.4 45.7 31.4 57.3G 18.7M 41.0
+RMP-SAM(Ours) R18 39.9 78.6 43.3 34.8 52.7 38.7 60.5G 22.8M 40.3
+Mask2Former R50 42.9 79.8 47.6 35.6 58.0 42.1 153G 45.2M 26.6
+MaskFormer R50 37.4 77.3 41.2 31.8 58.8 40.0 143.0G 45.2M 34.3
+kMaX-DeepLab R50 36.9 75.7 42.7 28.4 20.1 26.4 280.0G 51.4M 14.9
+YOSO R50 37.3 76.9 42.7 29.2 49.2 40.7 119.0G 45.1M 36.3
+RMP-SAM(Ours) R50 46.9 80.8 51.6 39.8 57.9 46.2 123.0G 47.2M 35.1
+MaskFormer TopFormer 31.6 75.9 33.5 28.6 56.1 38.2 60.0G 12.8M 29.9
+YOSO TopFormer 31.0 75.9 33.0 26.9 45.1 27.9 36.8G 12.9M 31.1
+RMP-SAM(Ours) TopFormer 34.6 77.0 37.8 29.8 53.3 41.7 40.1G 15.0M 30.7
 4.1 Main Results
@@
 We benchmark recent state-of-the-art methods using the same training and test settings.
-From the table, our proposed RAP-SAM achieves the best speed and accuracy trade-off on all three different visual segmentation tasks, under the various backbones.
-Mask2Former [10] achieves similar or partially stronger results than our method.
-However, their speed is still limited, which is 7-8 FPS slower than our methods.
-YOSO [25] runs as fast as our method.
+From the table, our proposed RMP-SAM achieves the best speed and accuracy trade-off on all three visual segmentation tasks under the various backbones.
+Mask2Former (Cheng et al., 2022) achieves similar or partially stronger results than our method.
+However, their speed is still limited, and they are 7-8 FPS slower than our methods.
+YOSO (Hu et al., 2023) runs as fast as our method.
 However, the performance gap is still significant.
-Thus, our proposed RAP-SAM is a simple yet effective baseline for real-time and all-purpose settings.
-Promptable Results Compared with SAM.
+Thus, our proposed RMP-SAM is a simple yet effective baseline for real-time and Multi-Purpose settings.
+Compare with SAM-like Methods.
+Tab.
+4 provides a detailed comparison between our models and previous SAM-like models.
+We report the results of mask AP of the COCO validation set, the parameter count, and FLOPs of each model.
+We first generate bounding boxes for each image using Mask R-CNN with ResNet50.
+Then, we send these bounding boxes to SAM-like models as box prompts.
+Our method is co-trained using SAM data and demonstrates even better results than most previous efficient SAM-like methods.
+It has strength in computational efficiency, with an additional function for video instance segmentation and panoptic segmentation.
+Comparison with Specific Design Models on VIP-Seg.
 In Tab.
-4, we verify our trained model as a promptable segmenter, like SAM.
-We compare our method with SAM-Huge, which is much larger.
-Despite there being a gap, our method can run fast and with much less GFlops.
-Comparison with Specific Design Models on VIP-Seg.
-We also verify the effectiveness of RAP-SAM on a more challenging video segmentation task, video panoptic segmentation on the VIP-Seg dataset.
-Compared with recent works [46, 61], RAP-SAM also archives the best speed and accuracy trade-off.
-Table 4: Comparison with SAM [34] as a promptable segmenter on various detectors and instance segmenters.
-Method Detectors mAP AP50 AP75 Parameters Flops
-SAM-Huge Mask-RCNN (R50) 35.6 54.9 38.4 641M 3000G
-Ours Mask-RCNN (R50) 26.7 43.0 27.9 47M 188G
-SAM-Huge Mask-RCNN (X-101-64x4d) 39.0 60.5 42.1 641M 3000G
-Ours Mask-RCNN (X-101-64x4d) 28.5 46.4 29.9 47M 188G
+5(b)(a), we verify the effectiveness of RMP-SAM on a more challenging video segmentation task, video panoptic segmentation (VPS) on the VIP-Seg dataset.
+Compared with recent works (Li et al., 2023e; Miao et al., 2022), RMP-SAM also archives the best speed and accuracy trade-off, despite RMP-SAM is not specifically designed for VPS.
+Table 4: Comparison with SAM-like methods on COCO instance segmentation dataset.
+We first generate bounding boxes using Faster R-CNN, which are used as visual prompts for segmentation.
+We report the results of mask AP of the COCO validation set, the parameters, and the FLOPs.
+Method mAP #Param. FLOPs(G) PS VIS Interactive
+SAM-H (Kirillov et al., 2023) 35.6 641M 3000
+TinySAM (Shu et al., 2023) 32.5 10M 42
+MobileSAM (Zhang et al., 2023a) 32.1 10M 42
+EdgeSAM (Zhou et al., 2023) 31.9 10M 22
+EfficientSAM (Xiong et al., 2023) 32.4 10M -
+FastSAM (Zhao et al., 2023) 34.3 68M 344
+RMP-SAM 33.2 47M 35
 Figure 3: Meta-architecture exploration.
@@
 Best viewed in color and zoom in.
-Table 5: Comparison with real-time video segmentation method on VIP-Seg validation set.
-Method backbone $\rm VPQ^{1}$ $\rm VPQ^{2}$ $\rm VPQ^{4}$ $\rm VPQ^{6}$ VPQ STQ FPS
-Clip-PanoFCN [61] R50 24.3 23.5 22.4 21.6 22.9 31.5 8
-Video K-Net [47] R50 29.5 26.5 24.5 23.7 26.1 33.1 10
-Tube-Link [46] STDCv1 32.1 31.3 30.1 29.1 30.6 32.0 14
-Tube-Link [46] STDCv2 33.2 31.8 30.6 29.6 31.4 32.8 12
-Ours R18 34.2 32.1 30.8 30.2 32.5 33.7 30
-Table 6: Comparison with real-time panoptic segmentation method on ADE-20k validation set.
-Method Backbone PQ $\mathrm{PQ^{th}}$ $\mathrm{PQ^{st}}$ Parameters Flops
-PanSegFormer [49] R50 36.4 35.3 38.6 51M 214G
-MaskFormer [11] R50 34.7 32.2 39.7 45.2M 181G
-Mask2Former [10] R50 39.7 39.0 40.9 45.2M 226G
-YOSO [25] R50 38.0 37.3 39.4 45.1M 176G
-RAP-SAM R50 38.3 36.6 39.7 47.2M 179G
-Comparison with Specific Design Models on ADE-20k.
-We further transfer RAP-SAM on ADE-20k datasets.
-Since our method is real-time model, we only list several representative works for reference.
-Again, compared with recent work [25], our method still achieves stronger results.
-5.2 Ablation Study and Visual Analysis
-Ablation on Shared Decoder Design.
-In Tab.
-7, we find using simple pooling-based dynamic convolution performs well under real-time settings, where we adopt R18 as the backbone and keep the adapter unchanged.
-In particular, there is no difference in panoptic segmentation results and a slight gap for interactive segmentation.
-However, per-pixel cross-attention introduces extra computation costs.
-Ablation on Meta-Architecture Design.
-In Tab.
-8, we further explore the meta-architecture design, as shown in Fig.
-3.
-We use R50 as the backbone and keep the adapter unchanged.
-From the table, we find using a shared decoder architecture (in Tab.
-7) achieves the best parameter and performance trade-off.
-These findings may conflict with recent multi-task works [42] since the decoupled decoders may have more capacity.
-We argue that since our feature extractors are weaker than those works and the representation power is far from the heavier backbone such as Swin [52] or ViT-large [13], adding more capacity in the decoder may not boost the final performance.
-Table 7: Ablation on shared decoder design.
-Setting COCO-Panoptic COCO-SAM FPS
-Per-Pixel Cross-Attention [10] 45.0 55.3 28.0
-Pooling + Dynamic Convolution [68] 43.8 55.7 36.2
-Pooling + Dynamic Convolution with Gate [93] 44.6 56.7 34.5
-Table 8: Ablation on meta-architecture.
+Table 5: Compared with specifically designed segmentation models.
+(a) Comparison with video segmentation method on VIP-Seg validation set.
+Method backbone VPQ STQ FPS
+Clip-PanoFCN R50 22.9 31.5 8
+Video K-Net R50 26.1 33.1 10
+Tube-Link STDCv1 30.6 32.0 14
+Tube-Link STDCv2 31.4 32.8 12
+Ours R18 32.5 33.7 30
+(b) Comparison with panoptic segmentation method on ADE-20k validation set.
+Method backbone PQ Parameters FLOPs
+PanSegFormer R50 36.4 51M 214G
+MaskFormer R50 34.7 45.2M 181G
+Mask2Former R50 39.7 45.2M 226G
+YOSO R50 38.0 45.1M 176G
+RMP-SAM R50 38.3 47.2M 179G
+Figure 4: The visualization results of YouTube-VIS 2019 and COCO datasets.
+The first two rows visualize five frames of inputs.
+The same instances are in the same color.
+The third row shows the interactive segmentation results with a single-point prompt (green color).
+The last row shows the panoptic segmentation results.
+Table 6: Ablation study on RMP-SAM’s component design.
+(a) Ablation on meta-architecture.
 Setting COCO-Panoptic COCO-SAM Parameters
@@
 Fig. 3 (d) 45.2 56.2 54.6M
-Table 9: Ablation on adapter design. CA: cross-attention. DC: dynamic convolution.
+(b) Ablation on joint co-training. a: COCO-Panoptic. b : YT-VIS 2019. c: COCO-SAM.
+Setting COCO-Panoptic YT-VIS 2019 COCO-SAM
+a 36.6 - -
+b - 21.5 -
+a + b 36.2 36.0 50.7
+a + b + c 35.7 35.3 50.9
+(c) Ablation on shared decoder design. DC: dynamic convolution. DCG: dynamic convolution with gating.
+Settings COCO-Panoptic COCO-SAM
+Per-Pixel Cross-Attention 45.0 55.3
+Pooling + DC 43.8 55.7
+Pooling + DCG 44.6 56.7
+(d) Ablation on adapter design. CA: cross-attention. DC: dynamic convolution.
 $\mathrm{A_{obj}}$ $\mathrm{A_{prompt}}$ COCO-Panoptic COCO-SAM
@@
 CA DC 44.5 56.2
-Table 10: Ablation on joint co-training. (a), COCO-Panoptic. (b), YT-VIS 2019. (c), COCO-SAM.
-Setting COCO-Panoptic YT-VIS 2019 COCO-SAM
-a 36.6 - -
-b - 21.5 -
-a + b 36.2 36.0 50.7
-a + b + c 35.7 35.3 50.9
-Figure 4: The visualization results of our model on the YouTube-VIS 2019 dataset and the COCO dataset.
-The first three rows visualize five frames of inputs.
-The same instances are shown with the same color.
-The fourth row shows the interactive segmentation results with a single-point prompt (shown in green color).
-The last row shows the panoptic segmentation results.
+Comparison with Specific Design Models on ADE-20k.
+In Tab.
+5(b)(b), we further transfer RMP-SAM on ADE-20k datasets.
+Since our method is a real-time model, we only list several representative works for reference.
+Our method still achieves stronger results than recent work (Hu et al., 2023).
+4.2 Ablation Study and Visual Analysis
+Ablation on Meta-Architecture Design.
+In Tab.
+6(d)(a), we further explore the meta-architecture design, as shown in Fig.
+3.
+We use R50 as the backbone and keep the adapter unchanged.
+The table shows that a shared decoder architecture achieves the best parameter and performance trade-off.
+These findings may conflict with recent multi-task work (Li et al., 2022c) since the decoupled decoders may have more capacity.
+We argue that since our feature extractors are weaker than those works and the representation power is far from the heavier backbone such as Swin (Liu et al., 2021) or ViT-large (Dosovitskiy et al., 2021), adding more capacity in the decoder may not boost the final performance.
+Effectiveness of Joint Co-training.
+We further explore the impact of each dataset in Tab.
+6(d)(b).
+We find joint co-training with image and video data leads to better performance for video instance segmentation but reduces the performance of panoptic segmentation.
+Adding COCO-SAM training leads to a few performance drops.
+This is because the learning targets are different for object queries and prompt queries.
+Ablation on Shared Decoder Design.
+Tab.
+6(d)(c) shows using simple pooling-based dynamic convolution performs well under real-time settings, where we adopt ResNet18 as the backbone and keep the adapter unchanged.
+In particular, there is no difference in panoptic segmentation results and a slight gap for interactive segmentation.
+However, per-pixel cross-attention introduces extra computation costs.
 Ablation on Adapter Design.
 Then, we explore the adapter design in Tab.
-9.
+6(d)(d).
 In particular, we use a pre-trained model without adapters for initialization to see the effect of different adapters.
-In particular, we find using asymmetric adapters works well for balanced results for object queries and prompt queries, since the goals of the two queries are different.
+In particular, we find that using asymmetric adapters works well for balanced results for object queries and prompt queries since the goals of the two queries are different.
 The former needs the spatio-temporal and scene-level context, while the latter only cares about the region context with the input position as the guidance.
-Effectiveness of Joint Co-training.
-We further explore the effect of each dataset in Tab.
-10.
-We find joint co-training with image and video data leads to better performance for video instance segmentation.
-Adding COCO-SAM training leads to a few performance drops.
-This is because the learning targets are different for object queries and prompt queries.
-Visualization.
+Our adapter can also work with other architecture, as shown in the appendix (Sec.
+7.2).
+Result Visualization.
 In Fig.
-4, we give the three different visualization results on the YouTube-VIS 2019 dataset and the COCO dataset.
+4, we give the three different visualization results on the YouTube-VIS 2019 and COCO datasets.
 We generate these visualization results using only a single model thanks to the joint co-training strategy.
@@
 4 show the video instance segmentation.
-The fourth row shows the interactive segmentation results with a single-point prompt (shown in green color), and the last row shows the panoptic results.
+The fourth row shows the interactive segmentation results with a single-point prompt (shown in green color), and the last row shows the results of panoptic segmentation.
+5 Related Work
+Universal Segmentation.
+Prior research has focused on designing segmentation models tailored for specific tasks.
+Recent developments in this field (Dosovitskiy et al., 2021; Liu et al., 2021; Carion et al., 2020; Zhang et al., 2021; Cheng et al., 2022; Yu et al., 2022; Li et al., 2021b; 2023c; Gu et al., 2023; Yuan et al., 2022; Li et al., 2022c; 2023d; 2024b; Yuan et al., 2024; Yang et al., 2021b; 2024) have shifted towards employing mask classification architectures, coupled with an end-to-end set prediction objective, to facilitate universal segmentation.
+This approach has outperformed specialized models (Chen et al., 2019; Kirillov et al., 2019; He et al., 2017; Li et al., 2021a; Huang et al., 2019; Yuan et al., 2020; Zhao et al., 2017; Chen et al., 2018; Li et al., 2022b; 2020a; Chen et al., 2017) across various segmentation tasks, including those involving images, videos, and point clouds (Kim et al., 2022; Li et al., 2022d; 2023e; Cheng et al., 2021a; Xu et al., 2022).
+Specifically, Mask2Former (Cheng et al., 2022) uses a masked-attention mechanism and surpasses the performance of prior specific image segmentation models.
+Similarly, Tube-Link (Li et al., 2023e) outperforms earlier models (Li et al., 2022d; b; Liu et al., 2020; Shin et al., 2023; Goel et al., 2021) tailored for specific video segmentation tasks.
+Semantic-SAM (Li et al., 2023a), built on Mask-DINO (Li et al., 2023b), employs a hybrid query design to facilitate panoptic and interactive segmentation within a single framework.
+However, despite the strong performance of these works, their complexity, stemming from the heavier encoder or the cascaded decoder designs, poses significant challenges for integration into actual products.
+Slow processing speeds and substantial parameter overhead primarily hinder these.
+Efficient Model Design.
+This research direction primarily concentrates on the development of efficient CNNs (Howard et al., 2017; Sandler et al., 2018; Howard et al., 2019; Iandola et al., 2016; Zhang et al., 2018; Ma et al., 2018; Han et al., 2020), transformers (Maaz et al., 2022; Pan et al., 2022; Mehta & Rastegari, 2022), and hybrid architectures (Zhang et al., 2023b; Ma et al., 2022; Chen et al., 2022b; Li et al., 2022a; Zhou et al., 2022), aimed at advancing visual representation learning.
+Most research investigates backbone design within the constraints of real-time processing and parameter efficiency.
+These studies are orthogonal to our work, which primarily explores the impact of data, task, and decoder design.
+Additionally, our work employs efficient models as encoders and presents detailed benchmarks and analyses of the encoder’s effects.
+Efficient Segmentation.
+Previous studies (Zhao et al., 2018; Li et al., 2020b; Yu et al., 2018; Hu et al., 2023; Hong et al., 2021; Wan et al., 2023; Yu et al., 2021; Mehta et al., 2018) on efficient segmentation have predominantly concentrated on closed-set and specific domains.
+In particular, a significant portion of this research (Li et al., 2020b; Hu et al., 2023; Mehta et al., 2019) is dedicated to driving scenarios.
+Recently, various studies (Zhang et al., 2022; Wan et al., 2023; Zhou et al., 2023) have developed efficient segmentation techniques that facilitate model execution on mobile devices.
+Mobile SAM (Zhang et al., 2023a) introduces a streamlined encoder distillation method.
+Fast SAM (Zhao et al., 2023) employs a single-stage instance segmentation framework that directly decodes class-agnostic masks.
+Recently, multiple studies have been conducted on efficient panoptic segmentation (Sun et al., 2023; Hu et al., 2023; Mohan & Valada, 2021) and rapid video instance segmentation (Yang et al., 2021a; Zhang et al., 2024).
+However, these real-time segmentation methods are limited to specific tasks.
+We contend that a multi-purpose model capable of real-time performance would have wide-ranging applications in editing, tracking, and segmentation functionalities within various products.
+Interactive Segmentation.
+Interactive segmentation (Kirillov et al., 2023; Chen et al., 2022a; Sofiiuk et al., 2022; Lin et al., 2020; 2022) distinguishes objects by actively incorporating user inputs.
+Recently, SAM (Kirillov et al., 2023) has shown remarkable segmentation outcomes by leveraging multiple user interaction inputs and training on SA-1B data.
+Inspired by SAM, several studies (Zhang et al., 2023a; Zhou et al., 2023) have concentrated on employing a lighter image encoder to adapt SAM for real-time scenarios.
 6 Conclusion
-In this work, we explore a challenging real-time setting, all-purpose segmentation, to transfer more segmentation tasks into one model.
-To solve this problem, we introduce RAP-SAM, a real-time segmentation model, that can segment and track objects in image, video, and interactive modes.
-We design a simple shared decoder with two key designs: shared dynamic convolution and asymmetric adapters, which lead to the best trade-off between speed and accuracy on three segmentation tasks.
-We hope our proposed benchmark inspires future research on this new setting.
-Overview.
-In this supplementary, we first present more detailed results to support both the benchmark and our proposed RAP-SAM in Sec.
-A.
-Then, in Sec.
-B, we present a more detailed visual comparison using our model.
-Finally, in Sec.
-C, we discuss several failure cases and potential future work.
-Appendix A More Experiment Results
-Benchmark More Methods.
-In Tab.
-11, we benchmark more models and backbones [93, 57, 58] as the supplementary to main paper.
-Again, for various backbones and different heads, our model can achieve better or comparable results across three different tasks, while running in real-time.
-In particular, compared with the dynamic kernel-based method, K-Net, our approach achieves better performance on the R18 and R50 backbones and comparable results on the lightweight backbones.
-Scale Up Joint Co-training.
-In the bottom of Tab.
-11, we further scale up our model on a large backbone, ConvNeXt large [53].
-From the table, we find there are significant gains in all settings while still running with 8 FPS.
-This means our proposed all-purposed segmentation still has room to balance the accuracy and inference speed.
-Table 11: Benchmark More Models on Real-Time All-Purpose Segmentation. The GFlops are obtained with $1333\times 800$ inputs.
-Method Backbone COCO-Panoptic COCO-SAM YouTube-VIS 2019 Flops Parameters FPS
-PQ SQ $\mathrm{PQ_{th}}$ $\mathrm{PQ_{st}}$ mIoU mAP
-YOSO [25] EdgeNexT [57] 31.8 74.5 35.6 26.1 47.4 35.4 40G 15.3M 28.6
-K-Net [93] R18 [21] 33.1 75.7 36.8 27.4 53.6 34.8 124G 25.6M 30.4
-K-Net [93] R50 [21] 39.2 78.2 43.9 31.6 56.3 41.7 171G 38.5M 20.9
-K-Net [93] EdgeNexT [57] 38.0 76.7 42.5 31.4 55.7 44.0 108G 19.5M 30.5
-Mask2Former [10] EdgeNexT [57] 39.8 78.8 43.6 34.0 54.2 - 73G 12.0M 25.6
-RAP-SAM R18 39.9 78.6 43.3 34.8 52.7 38.7 60.5G 22.8M 40.3
-RAP-SAM R50 46.9 80.8 51.6 39.8 57.9 46.2 123.0G 47.2M 35.1
-RAP-SAM EdgeNexT [57] 38.0 77.7 42.1 31.9 54.8 46.3 44G 14.3M 31.5
-RAP-SAM ConvNeXt-L [53] 52.3 82.5 58.6 42.8 61.1 55.6 700G 239M 8.6
-Explore adapter Design on More Methods.
-We explore the effectiveness of our adapter design on Mask2Former [10] and add the adapter module to the last decoder layer of Mask2Former.
-As shown in Tab.
-12, Mask2Former with adapter will improve 1.1 mIoU score compared with origin Mask2Former for interactive segmentation and only drop 0.3 PQ for panoptic segmentation.
-This result indicates that the adapter module not only applies to our method but can also be generalized to other methods
-Table 12: Explore adapter Design on More Methods.
-Method Backbone PQ mIoU Mask2Former R50 43.2 57.0 Mask2Former+adapter R50 43.0 58.1
-More Implementation Details.
-We implement our proposed method and benchmark methods using MMDetection toolbox [5], and we train all of these methods in the same environment.
-For kMaX-DeepLab [85], we excluded the auxiliary semantic segmentation loss, instance discrimination loss, and preserving PQ-style loss alignment with Mask2former [10] for a fair comparison.
-We utilize the AdamW [54] optimizer with a weight decay of 0.05, and the learning rate is set to 1e-4 for all methods.
-We warm up the learning rate in the first 500 iterations using a linearly increased strategy and decay the learning rate at 8 and 11 epochs by a factor of 10.
-For data augmentation, we use the large-scale jittering (LSJ) augmentation [16] with a random scale sampled from the range of 0.1 to 2.0 and followed by a fixed size crop to 1024×1024.
-More Inference Details.
-For video instance segmentation, we adopt simple object query matching [47, 26] for all methods to keep the fair comparison.
-For interactive segmentation, we mainly use point prompts, which is more challenging than box prompts.
-For panoptic segmentation, all hyperparameters are the same.
-The FPS is obtained on one A100 card for both the main paper and supplementary.
-Appendix B Visual Comparison
-Comparison on Youtube-VIS dataset.
-In Fig.
-5, we compare our RAP-SAM (right) with a strong baseline Mask2Former (left) with the same ResNet50 backbone.
-Our methods achieve more consistent tracking and segmentation results in these examples.
-Figure 5: When compared to the Mask2Former(right) on the YouTube-VIS 2019 dataset, our RAP-SAM(left) demonstrates superior performance in recognizing and segmenting objects in certain scenarios, and it also holds an advantage in detailing edges.
-Comparison on COCO Panoptic Segmentation dataset.
-In Fig.
-6, We demonstrate that our RAP-SAM model surpasses K-Net in panoptic segmentation tasks.
-RAP-SAM exhibits enhanced accuracy in processing complex scenes and maintaining fine details, especially in segmenting object edges and managing overlapping objects.
-Figure 6: When compared with K-Net (bottom) on the COCO dataset, our RAP-SAM (top) model shows a significant advantage, achieving more accurate segmentation of foreground and background, with edges that are smoother.
-More Interactive Segmentation Results.
-In Fig.
-7, We present further visualizations of interactive segmentation, showcasing segmentation outcomes for ’things’ and ’stuff’.
-In these visualizations, green dots signify manually specified prompts.
-Our model exhibits precise segmentation abilities for ’things’ and ’stuff’.
-Notably, even when prompt points are situated on object boundaries, our model accurately identifies and segments the object.
-Figure 7: Here are additional results of interactive segmentation.
-For point prompts or box prompts, RAP-SAM segments out the corresponding object based on the spatial location of the prompt.
-We have provided both segmentation results for ’things’ and ’stuff’, demonstrating that RAP-SAM is capable of segmenting both while maintaining strong performance.
-Appendix C Challenges and Future Work
-Failure Cases Analysis.
-We also analyze failure cases in video/image segmentation and identify two typical failure modes of RAP-SAM.
-Firstly, as illustrated in Fig.
-8, when faced with multiple objects with high overlap, RAP-SAM struggles to fully recognize them, resulting in the omission of several masks.
-Secondly, in crowded scenarios, RAP-SAM faces challenges in recognizing and segmenting all instances with a limited number of instance kernels.
-Figure 8: Failure modes of RAP-SAM in image/video segmentation.
-Future Work Discussion.
-There are several directions to explore for real-time all-purpose segmentation.
-The first direction is to achieve better-balanced performance on image, video, and interactive segmentation (since there is still a lot of room for performance, as shown in the last row of Tab.
-11).
-The second direction is to speed up the model and deploy the model on the edge device, such as iPhone.
-The third direction is to explore different knowledge distillation approaches to transfer the vision foundation model in real-time all-purpose models.
-Moreover, more visual prompts can be explored, such as mask prompts or box prompts.
-Our method currently only supports point prompts.
-This will be our future work.
+In this work, we explore a new challenging setting, real-time multi-purpose segmentation, to explore multi-task segmentation in one efficient model.
+To achieve this, we first benchmark existing image segmentation approaches by extending their query for video and interactive segmentation.
+However, no work can achieve the best speed and accuracy trade-off on all three sub-tasks.
+To solve this problem, we introduce RMP-SAM, a real-time segmentation model.
+It has a simple shared decoder with two key designs: shared dynamic convolution and asymmetric adapters.
+These designs lead to the best trade-off between speed and accuracy on three segmentation tasks.
+The proposed RMP-SAM model is versatile and effective for various tasks, including image segmentation, video segmentation, and iterative segmentation.
+To our knowledge, no real-time model can achieve this goal.
+The proposed benchmark and RMP-SAM baseline can facilitate future research in this new setting.
+Acknowledgments
+This work is supported by the National Key Research and Development Program of China (No.2023YFC3807603).